### v1

# PyFi: Toward Pyramid-like Financial Image Understanding for VLMs via Adversarial Agents

Abstract
This paper proposes PyFi, a novel framework for pyramid-like financial image understanding that enables vision language models (VLMs) to reason through question chains in a progressive, simple-to-complex manner. At the core of PyFi is PyFi-600K, a dataset comprising 600K financial question-answer pairs organized into a reasoning pyramid: questions at the base require only basic perception, while those toward the apex demand increasing levels of capability in financial visual understanding and expertise. This data is scalable because it is synthesized without human annotations, using PyFi-adv, a multi-agent adversarial mechanism under the Monte Carlo Tree Search (MCTS) paradigm, in which, for each image, a challenger agent competes with a solver agent by generating question chains that progressively probe deeper capability levels in financial visual reasoning. Leveraging this dataset, we present fine-grained, hierarchical, and comprehensive evaluations of advanced VLMs in the financial domain. Moreover, fine-tuning Qwen2.5-VL-3B and Qwen2.5-VL-7B on the pyramid-structured question chains enables these models to answer complex financial questions by decomposing them into sub-questions with gradually increasing reasoning demands, yielding average accuracy improvements of 19.52% and 8.06%, respectively, on the dataset. All resources of code, dataset and models are available at: https://github.com/AgenticFinLab/PyFi.
PyFi: Toward Pyramid-like Financial Image Understanding for VLMs via Adversarial Agents
Yuqun Zhang1∗     Yuxuan Zhao12∗     Sijia Chen1† 1 The Hong Kong University of Science and Technology (Guangzhou) 2 Yantai Research Institute, Harbin Engineering University research@yuqunzhang.com, yx.zhao129@gmail.com, sijiachen@hkust-gz.edu.cn
footnotetext: Equal contribution.footnotetext: Corresponding author.
1 Introduction
Vision language models (VLMs) are capable of step-by-step reasoning, presented as Chain-of-Thought (CoT) CoT-neurips22, for image understanding in question answering (Q&A) across various scientific domains (MMStar-neurips24; MMMU-cvpr24; Visco-cvpr25), driven by the development of domain-specific benchmarks (VisNumBench-iccv25; Mvmath-cvpr25; MMMU-cvpr24; Codevision-arxiv25), step-wise annotations (MathShepherd-arxiv23; o1-coder-arxiv24; mutualreasoning-arxiv24), and domain-tailored VLMs (FinGPT-ijcai23; open-finllms-arxiv24; finlmm-r1-arxiv25). In contrast, advancements in the financial domain remain limited. This is because, in finance image understanding, the creation of CoT reasoning datasets, which are highly necessary for the evaluation and fine-tuning of VLMs, requires a high level of expertise and domain knowledge and is therefore scarce.
Figure 1: Overview of the PyFi-600K dataset within our PyFi framework. The dataset is structured as a pyramid comprising 6 capability levels, $1$ financial image types, and 17 financial themes.
To address these issues, previous literature (FinQA-emnlp21) collects Q&A samples from financial documents manually, a process that is prohibitively time-consuming and difficult to scale. As a result, subsequent works propose datasets by merging extensive existing financial datasets (FLUE-emnlp22; Finben-neurips24). These repurposed datasets may introduce accumuated biases and inconsistencies inherited from the original sources. More recently, benchmarks such as (Mme-finance-mm25; FCMR-acl25; FinMME-acl25; BizFinBench-arxiv25; CFBenchmark-MM-arxiv25; VisFinEval-arxiv25) leverage trained VLMs to generate Q&A pairs augmented with CoTs (FAMMA-arxiv25). However, these samples may contain factual errors or hallucinations, which can distort VLM evaluation and fine-tuning and thereby undermine their reliability in real-world financial applications, where precision and expert-level understanding are essential.
Therefore, this paper proposes a framework named PyFi, which supports comprehensive evaluation and advanced fine-tuning of VLMs through a pyramid-like hierarchy for financial image understanding. Specifically, PyFi includes a synthesized dataset, PyFi-600K shown in Figure 1, comprising 600 thousand (K) samples with step-wise annotations that provide expert-level guidance on how to progressively solve complex financial problems across $17$ categories. The construction of this dataset is enabled by PyFi-adv, a multi-agent adversarial mechanism designed to automatically synthesize and refine samples using expert-level financial knowledge.
PyFi-600K offers three key benefits for the interpretability and fine-tuning of VLMs in finance. First, the samples are organized in a pyramid structure, where the required image understanding capability increases progressively — from basic visual perception at the bottom level to complex financial decision-making at the top. This hierarchy, with capability levels ranging from one to six, enables comprehensive evaluation of VLMs across multiple aspects and supports ranking the capabilities required by VLMs in different financial themes. Second, samples across different levels are interconnected through question chains, demonstrating how questions at higher levels can be addressed step-by-step by leveraging information and reasoning from lower-level Q&A samples. Finally, it supports process supervision (PRM800K-iclr23), as each sample is assigned a reward score that reflects the correctness of its reasoning process in solving the financial problem.
The construction of such a scalable dataset is enabled by our proposed PyFi-adv, a multi-agent adversarial mechanism in which a challenger agent progressively synthesizes increasingly difficult questions, while a solver agent generates answers by reasoning with financial knowledge. The two agents compete within a Monte Carlo Tree Search (MCTS) framework: the challenger begins with simple questions at level 1, either selecting existing ones or generating new ones, and presents them to the solver for response. Based on the answer, the challenger then formulates a more challenging question in the next round. This iterative process continues until the information accumulated suffices to address the decision-making task at level 6.
Our extensive experiments with PyFi-600K reveal that 15 well-known VLMs struggle to answer complex financial questions, with average accuracy dropping from 71.80% at level 1 to 32.95% at level 6. With question chains, we observe that addressing a Level-6 financial decision-making question typically requires correctly answering an average of $10.48$ sub-questions from lower levels. In particular, our results highlight that errors in calculation analysis are a primary cause of VLMs’ failure in final financial decision-making. More importantly, fine-tuning Qwen-VL models on PyFi-600K using question chains improves accuracy by an average of 13.79%, with gains reaching up to 19.52% in the best case.
2 Related Work
Financial Benchmark. The evaluation of Vision-Language Models (VLMs) in the finance domain has seen remarkable progress, from early text-based benchmarks (FinQA-emnlp21; FLUE-emnlp22; Finben-neurips24) to more recent efforts targeting financial image understanding (FCMR-acl25; FinTMMBench-mm25; FinChartBench-arxiv25; CFBenchmark-MM-arxiv25; VisFinEval-arxiv25; Fin-Fact-mm25; FinRAGBench-V-arxiv25; Mme-finance-mm25; FinMME-acl25; FAMMA-arxiv25; FinMR-iclr25). However, the question-answering datasets of most benchmarks, such as FinMME (Mme-finance-mm25), rely on human annotation, making them costly and difficult to scale. Besides, they fail to support comprehensive assessment of VLMs in the financial domain. However, we introduce an automatically synthesized dataset for the fine-grained, hierarchical evaluation of financial visual understanding across diverse image types and financial themes.
Financial Large Models. Inspired by the success of large models (LMs) in general domains, recent research has increasingly focused on developing specialized financial LMs to address tasks requiring deep domain expertise. Previous efforts primarily centered on text analysis using financial LMs (FinBERT-car23; Bloomberggpt-arxiv23; FinGPT-ijcai23; PIXIU-neurips23; Finr1-arxiv25). Recent financial VLMs (Finvis-gpt-ijcai23; FinTral-acl24; open-finllms-arxiv24; FinTab-LLaVA-kdd25) integrate visual inputs to facilitate step-by-step reasoning, thereby enabling broader financial decision-making. However, these models lack the expertise and interpretability needed for challenging financial themes, a limitation stemming from the absence of detailed, step-wise problem-solving processes grounded in finance-specific knowledge. In contrast, our dataset contains questions that illustrate how to address financial problems through a progression from simple to complex image understanding and from basic to increasingly deep financial knowledge.
Multi-Agent LM Framework. Recent advances in LMs have enabled Agentic AI to excel in real-world applications, with multiple LM-based agents now collaborating to perceive, learn, reason, and act, supporting intricate tasks such as simulation (Multi-agent-simulation-arxiv25; Multi-agent-simulator-acl25), reasoning (BeTop-neurips24; AgenticReasoning-neurips25; MultiAgentDebate-icml24), trading (Hedgeagents-www25; FinCon-neurips24; MountainLion-arxiv25), and data synthesis (systematic-simulation-arxiv24; STORM-BORN-acl25; AgentSGEN-arxiv25). However, to our knowledge, no existing method in finance leverages a multi-agent adversarial approach to synthesize data. In contrast, inspired by Generative Adversarial Networks (GANs) gan-neurips14, we present the first method that employs adversarial agents to automatically synthesize hierarchical, fine-grained financial datasets.
3 Framework
Figure 2: Overview of PyFi-adv: a challenger agent competes with a solver agent under the MCTS paradigm to generate question chains that progressively probe deeper capability levels in financial visual reasoning.
3.1 Financial Dataset in Pyramid Structure
We release PyFi-600K, a large-scale finance dataset comprising 600 thousand (K) samples, each consisting of a question-answer pair, constructed within a pyramid structure that encompasses six levels of finance-specific visual cognition: Perception (PP), Data Extraction (DE), Calculation Analysis (CA), Pattern Recognition (PR), Logical Reasoning (LR), Decision Support (DS). More importantly, PyFi-600K introduces sample chains that demonstrate how to address higher-level pyramid questions by leveraging knowledge from lower-level samples.
Specifically, we denote each finance sample as $\bm{s}^{l}=\left(I,Q,A,r\right)$, where $r\in\left[0,1\right]$ is the reward reflecting its correctness, $I$ is the finance image to understand, $l\in\left[1,6\right]$ indicates the pyramid level to which the question $Q$ and the answer $A$ belong. The sample complexity degree is $C\left(\bm{s}^{l}\right)\in\left[1,5\right]$. And, as presented in right sub-figure of Figure 2, we define sample chain as $\bm{S}^{l}_{1\dots n}=\left[\bm{s}^{1}_{1},...,\bm{s}^{l}_{n}\right]$ where $n$ is the length and $n\geq l$ because samples within the chain can come from the same level. Considering the chain, the sequence of questions increases in levels or complexity degrees, as the answer for each one requires more insightful image understanding and financial expertise than the preceding ones. That is, for any $\bm{s}^{h}_{i}$ and $\bm{s}^{k}_{j}$ with $i,j\in[1,\dots,n]$, $h,k\in[1,\dots,l]$, $i<j$, we have that $h\leq k$, and if $h=k$, then $C\left(\bm{s}^{h}_{i}\right)\leq C\left(\bm{s}^{k}_{j}\right)$. Subsequently, the reasoning process for answering the question in any sample within the chain depends on its predecessor samples. Specifically, when $\bm{s}^{l}_{j}$ is addressed by a VLM parameterized by $\bm{\theta}$, we have $p_{\bm{\theta}}\left(A|I,Q\right)<p_{\bm{\theta}}\left(A|I,Q,\left\{\bm{s}^{l}_{i}\right\}^{j}_{i=1}\right)$, meaning that the reasoning information and financial knowledge in answers from predecessor samples enable more reliable question answering for the current sample.
Therefore, we observe the following three benefits for the finance domain:
Hierarchical Cognition in the dataset embodies human-level problem solving processes, progressing from basic perception to complex financial decision-making. Samples at each level $l$ correspond to a specific cognitive capability, which builds upon those from lower levels i=1..,l1i=1..,l-1 or complexities. Thus, evaluating VLMs on these hierarchically structured samples enables a comprehensive assessment of their abilities in the finance domain. That is, any financial reasoning performed for questions from higher levels, which require rare, expert-level knowledge, relies on foundational financial questions involving base cognitions such as CA. Besides, we can identify how VLMs perform, whether better or worse, at each cognition level during financial reasoning, thereby revealing which level is crucial for which financial questions and to what extent.
Progressive Logical Chain in the dataset is presented as the sample chain $\bm{S}^{l}_{1\dots n}$ shown in Figure 2, which demonstrates how a challenging financial question at level $l=6$ is solved by progressively addressing a sequence of simple-to-complex sub-questions. This aligns with financial decision-making logic, where each step builds upon the previous one to form a cumulative and logically coherent reasoning flow. Such a traceable chain enhances the interpretability and justifiability required in the finance domain. Moreover, VLMs fine-tuned on such chains are capable of solving complex financial problems by accumulating knowledge through first addressing a series of simpler problems.
Step-wise Reward Score $r$ of each sample reflects its reliability; thus, the scores along a sample chain support process supervision (PRM800K-iclr23) in reasoning with financial image understanding. Using these scores, we enable the training of process reward models that act as verifiers for any financial VLM.
3.2 Automatic Synthesization with Adversarial Agents
Existing financial datasets rely solely on human annotation, while automated methods from other domains, such as rStar-Math (rStar-Math-arxiv25) for mathematical reasoning, fail to meet the requirements of finance, where interpretability and reliable reasoning are essential, along with increasing complexity, domain expertise, and hierarchical cognition. We achieve this by proposing a multi-agent adversarial mechanism, referred to as PyFi-adv, which constructs question chains, as illustrated in Figure 2
Specifically, PyFi-adv contains a challenger agent that is supported by an off-and-shelf VLM, denoted as $\psi$, and similarly, a solver agent $\phi$. Inspired by the confrontation in the game of Go used in AlphaGo (AlphaGo-nature), our two agents compete within a MCTS framework to automatically synthesize samples.
Confrontation for Sample Synthesization, as shown by the Gray oval circle in Figure 2, is that when $\phi$ aims to correctly answer the given question, $\psi$ checks the answer to propose a more challenging follow-up question to hinder the solver. This is formulated as $Q_{i}\sim p_{\psi}\left(Q_{i}|I,{\bm{\mathsfit{P}}}_{q},\bm{S}^{l}_{1\dots i-1}\right)$ and thus $A_{i}\sim p_{\phi}(A_{i}|I,Q_{i},{\bm{\mathsfit{P}}}_{a},\bm{S}^{l}_{1\dots i-1})$, where ${\bm{\mathsfit{P}}}_{q}$ and ${\bm{\mathsfit{P}}}_{a}$ are the agent prompts.
Exploration for Diversity Guarantee aims to prevent the adversarial process from converging to a set of questions with identical content. At each step, a Bernoulli sampling is used to determine the $\psi$ whether to explore a new question path or exploit an existing one. For exploitation, the UCT formula is employed to select the most promising prior question.
Expansion for Hierarchical Questions is naturally achieved through the adversarial agents: once $\psi$ performs a question expansion, it generates a subsequent, more complex question based on the accumulated $\bm{S}^{l}_{1\dots i}$. Agent $\phi$ then produces an answer $A_{i+1}$, forming a new sample $\bm{s}^{l}_{i+1}$. Thus, a hierarchical question structure is built through repeated adversarial interactions that continue until either the top-level ($l=6$) question is addressed or a predefined depth limit is reached.
Backpropagation for Reward Assignment adjusts the success rate of each sample by backpropagating the comparison between the answer of the top-level question and the ground truth down to individual samples in the chain, as indicated by the orange dotted line in Figure 2. Upon convergence of the MCTS, the resulting success rate is used as the reward score, providing a correctness label that measures the reliability of the synthesized dataset.
4 Experiments
4.1 Setup
Table 1: Evaluation of 15 Pre-trained VLMs on PyFi-600K across $6$ capability levels, 55 complexity degrees. For simplicity, we only present the results from degrees 33, 44, and 55. The test set consists of $301$ samples, derived from an initial random selection of $1000$ instances followed by an additional filtering step. The full model names listed in the “Model” column are: InternVL3-38B, Claude-opus-4-1-20250805, DeepSeek-VL2, Hunyuan-Large-Vision, ERNIE-4.5-turbo-vl, Moonshot-V1-128k-Vision-Preview, Moonshot-V1-32k-Vision-Preview, Moonshot-V1-8k-Vision-Preview, Qwen3-VL-Plus, Qwen2.5-VL-72B-Instruct, Qwen2.5-VL-32B-Instruct, Qwen2.5-VL-7B-Instruct, and Qwen2.5-VL-3B-Instruct.
Model Overall PP DE CA PR LR DS 33 44 55 GPT-4.1 52.99 86.84 47.64 40.62 57.14 54.84 38.46 52.63 46.55 53.85 InternVL3 52.91 76.32 56.37 44.53 43.88 53.23 30.77 41.78 44.83 67.31 Claude-opus 64.70 80.92 62.74 66.02 58.16 72.58 32.69 66.12 55.60 73.08 GLM-4.5V 74.75 89.47 78.07 65.23 75.51 75.81 46.15 72.37 64.22 73.08 DS-VL2 45.18 76.32 44.58 36.33 42.35 42.74 19.23 34.54 42.24 46.15 Hunyuan-LV 59.72 90.79 58.96 36.72 63.27 69.35 51.92 50.00 57.76 80.77 ER-4.5-TVL 34.47 34.87 33.96 36.72 25.00 45.16 36.54 38.82 25.43 65.38 MV1-128k 54.57 80.92 61.08 39.06 41.84 58.87 38.46 46.71 43.97 57.69 MV1-32k 54.40 80.26 61.56 37.11 40.31 62.90 38.46 45.72 43.97 61.54 MV1-8k 54.90 80.26 61.79 39.84 42.35 58.06 38.46 48.03 43.97 55.77 Q3-VL-Plus 51.00 77.63 54.25 37.50 41.84 57.26 32.69 45.72 37.50 65.38 Q2.5-VL-72B 48.84 70.39 59.91 33.20 35.71 46.77 26.92 36.51 31.03 57.69 Q2.5-VL-32B 43.19 60.53 45.28 37.50 36.73 42.74 28.85 39.14 30.60 59.62 Q2.5-VL-7B 37.87 54.61 49.53 34.77 18.37 24.19 15.38 31.91 17.67 38.46 Q2.5-VL-3B 20.51 36.84 15.57 21.48 16.84 21.77 19.23 23.03 18.97 30.77
Datasets. PyFi-600K dataset consists of 62,660 sample chains with reward scores of 1. For evaluation, we proportionally sample 1,000 samples from each capability level to construct the test set across all cases. For supervised fine-tuning (SFT), we directly filter out samples for which more than 8 out of 15 VLMs produce the same answer even when the financial image is missing. This additional filtering step ensures the removal of samples that may have been memorized by VLMs during pre-training, a phenomenon referred to as data leakage.
Models. For evaluation, we use 15 VLMs as listed in Table 3. For SFT, we fine-tune Qwen2.5-VL-3B-Instruct and Qwen2.5-VL-7B-Instruct on sample chains to obtain PyFi-QwenVL-3B and PyFi-QwenVL-7B, respectively.
SFT. We fine-tuned the models on approximately 47K sample chains using two dataset variants: one with chain-of-thought (CoT) annotations (w/ CoT) and one without (w/o CoT). In the CoT variant, each question chain is converted into a CoT reasoning sequence, where each reasoning step corresponds to a sub-question and its answer. Thus, a chain with $n$ samples yields a CoT sequence with $n$ reasoning steps. In the non-CoT variant, only the question and answer from the final sample in the chain are used. We use the AdamW optimizer with a learning rate of $1.0\times 10^{-4}$, cosine learning rate scheduling, and a warmup ratio of 0.1. Training is conducted for one epoch with an effective batch size of 8. For parameter-efficient fine-tuning, we apply LoRA with full-module adaptation. All experiments are carried out on four NVIDIA RTX 5090 GPUs.
[…]
4.2 Main Results
Evaluations of $15$ Pre-trained VLMs are presented in Table 1. In summary, we observe that as the capability level and complexity of financial image understanding increase, the accuracy of VLMs gradually declines. Specifically, the average accuracy across all models decreases from 71.80% at level 1 (Perception) to 32.95% at level 6 (Decision Support). This indicates that while current VLMs can handle basic visual perception in the financial image understanding, they struggle with complex financial questions that require higher-level cognitive abilities. Notably, GLM-4.5V achieves the highest overall accuracy of 74.75%, demonstrating its strong performance across various capability levels. However, even this model’s accuracy drops to 46.15% at level 6, highlighting the challenges in financial decision-making. Most models, including GPT-4.1 and InternVL3-38B, achieve over 70% accuracy at level 1 but fall below 40% at level 6, resulting in a relatively low average overall accuracy of 50%.
Evaluations of Fine-tuned Qwen-VL models are presented in Figure 3. In summary, with question chains arranged in a pyramid manner, the models, especially the smaller ones, show a significant increase in accuracy and gain the ability to reason progressively, moving from low-level to high-level questions until they reach the solution, as shown in Figure 5. Specifically, both our PyFi- models exhibit substantial accuracy gains. PyFi-CoT related models improves the 3B model by 19.52% points and the 7B model by 8.06% points over their local baselines. The gain is notably larger for the smaller 3B model. Notably, on higher-level financial questions, such as Decision Support (DS), PyFi-CoT ones improve the 3B model by 9.62% and the 7B model by 23.08%. The substantially larger gain for the 7B VLM suggests that larger VLMs benefit more from question-chain fine-tuning in challenging financial scenarios. Notably, in Pattern Recognition (PR) and Logical Reasoning (LR), PyFi-QwenVL-3B even surpasses the accuracy of the much larger 72B VLM, as shown in Table 1. Building on the detailed comparison in Figure 5, we argue that sample-chain fine-tuning equips VLMs with the ability to solve complex financial problems by iteratively addressing a sequence of self-generated sub-questions of increasing capability levels, thereby significantly improving interpretability and reliability.
Figure 3: Comparison between Qwen2.5-VL models and ours ( PyFi-QwenVL-3B and PyFi-QwenVL-7B).
Figure 4: Evaluation of error proportions across capability levels leading to incorrect financial decisions. As illustrated in the right sub-figure, the sample chain enables error tracing: when Claude-Opus-4-1 fails to answer the Level 6 financial decision-making question, we can trace back to identify the specific lower-level sub-questions that were answered incorrectly, thereby pinpointing the exact level at which the error originated.
4.3 Main Benefits and Insights
Hierarchical Evaluation of VLMs in Table 1 and Figure 3 enables a fine-grained analysis of how — and to what extent — VLMs meet the varying capability requirements of financial image understanding across different levels of complexity. The phenomenon of decreasing accuracy as the level increases in Table 1 indicates that existing VLMs can only meet the most basic image understanding in finance and are difficult to be used for complex financial decision-making. For instance, while models like GLM-4.5V perform well on basic perception questions (level 1), their accuracy drops considerably on decision support questions (level 6). This trend is consistent across most models, suggesting that greater efforts are needed to enhance VLMs’ capabilities in the finance domain. Thus, the results in Figure 3 demonstrate the importance of our question-chain fine-tuning: Qwen models learn to make accurate financial decisions by first addressing a series of sub-questions in a progression from simple to complex, a process aligned with human cognition.
Tracing back from Decision Support to Lower-level Q&As, as illustrated in Figure 4, significantly improves the interpretability of VLMs in financial decision-making. In Figure 4, we first collect Level-6 samples that VLMs fail to answer correctly and analyze their question chains by evaluating VLMs on the questions at each lower level in the chain. In the finance domain, addressing a Level-6 financial decision-making question typically requires correctly answering sub-questions from lower levels. Thus, when an error occurs, as indicated by the red cross in the right sub-figure, PyFi-600K enables us to trace the problem-solving chain from DS to LR and eventually to PP, thereby diagnosing how an error in the DE leads to the final mistake. Such a tracing-back approach is crucial for understanding the limitations of VLMs in financial decision-making, as it reveals at which capability level a failure occurs. It is also essential for guiding how to adjust the reasoning chain to achieve accurate final decisions.
Identifying the Most Frequent Failure Capability Levels, as presented in the left part of Figure 4, reveals that the key to improving VLMs’ performance in the financial domain lies in strengthening their capabilities at the CA (Calculation Analysis) level. Specifically, the pyramid-like structure and the question chain of PyFi-600K facilitates detailed error analysis, allowing us to identify specific levels where VLMs struggle. That is, most VLMs, even the advanced Claude-opus-4-1-20250805 and Qwen3-VL-Plus, fails to answer more than 40% of the question from CA, revealing that the calculation analysis is a critical bottleneck for them. More importantly, by comparing the error proportions between six capability levels, we can prioritize which capabilities to enhance first. For instance, since CA has the highest error proportion, focusing on improving calculation analysis skills in VLMs could lead to the most significant overall performance gains in financial decision-making tasks. In addition, this comparison also shows that, in the finance domain, analyzing numerical calculations from images is one of the most critical capabilities for accurate decision-making, even more important than logical reasoning, aligning with conclusions drawn by real-world finance experts  [^1] We consulted 15 experts from financial quantification, market simulation, investment practice, and academia to arrive at this conclusion..
Figure 5: Qualitative analysis of level-wise COT by comparing the Qwen2.5-VL-3B-Instruct and PyFi-QwenVL-3B in the challenging financial decision-making.
Addressing Financial Decision-Making with Level-wise CoT, as depicted in Figure 5, is not only reliable but also highly interpretable. Specifically, we compare the responses of Qwen2.5-VL-3B-Instruct and PyFi-QwenVL-3B — the latter fine-tuned on question chains— when addressing the financial problem. The Qwen2.5-VL-3B-Instruct produces a short and disorganized reasoning process, reflecting its limited financial reasoning capability. In contrast, PyFi-QwenVL-3B follows a clear, structured, and logical reasoning flow, which is essential for expert-level performance in finance. Mirroring the pyramid structure of PyFi-600K and the progressively increasing capability level of questions in the question chains, our model first addresses basic PP questions, then proceeds to DE and CA, and finally tackles high-level LR and DS. This level-wise CoT not only improves answer reliability but also provides a transparent and traceable reasoning path that users can easily follow and verify, making it especially valuable in financial contexts where stakes are high. Moreover, experts can inspect this professional decision-making process and provide detailed and direct feedback to further refine the model.
Providing Process Supervision, in the form of a reward score for each sample in our PyFi-600K, enables research into training reliable verifier models, known as process reward models (PRMs), to facilitate checking the correctness of each step generated by financial VLMs. Even though, due to space limitations, this paper does not perform such training, we use the reward scores to filter out low-quality samples during both evaluation and fine-tuning, as discussed in the experimental setup.
5 Concluding Remarks
In this paper, we have proposed a framework named PyFi, containing the dataset PyFi-600K and the multi-agent adversarial mechanism PyFi-adv, to facilitate pyramid-like financial image understanding for VLMs. With the Q&A samples organized in a pyramid structure, PyFi-600K enables comprehensive evaluation of VLMs across $6$ capability levels, $17$ financial themes and $11$ image types. Our novel level-wise question chain supports the interpretability of the performance of VLMs in financial decision-making by 1) showing importance ranking of capability level in accurate decisions and 2) enabling a back-track of which errors in lower levels led to the decision failure. More importantly, with PyFi-adv, the dataset is scalable without human annotations, meaning the size and quality can be improved continuously. After performing extensive experiments on $15$ VLMs and especially fine-tuning on Qwen models with question chains, we have verified these benefits of the PyFi by showing detailed, hierarchical, and fine-grained results, exposing important insights such as Calculation Analysis (CA) is the most crucial for correct decision-making and fine-tuned small VLMs gain $19.52\%$ accuracy improvement.