### v2

# PyFi: Toward Pyramid-like Financial Image Understanding for VLMs via Adversarial Agents

Abstract
This paper proposes PyFi, a novel framework for pyramid-like financial image understanding that enables vision language models (VLMs) to reason through question chains in a progressive, simple-to-complex manner. At the core of PyFi is PyFi-600K, a dataset comprising 600K financial question-answer pairs organized into a reasoning pyramid: questions at the base require only basic perception, while those toward the apex demand increasing levels of capability in financial visual understanding and expertise. This data is scalable because it is synthesized without human annotations, using PyFi-adv, a multi-agent adversarial mechanism under the Monte Carlo Tree Search (MCTS) paradigm, in which, for each image, a challenger agent competes with a solver agent by generating question chains that progressively probe deeper capability levels in financial visual reasoning. Leveraging this dataset, we present fine-grained, hierarchical, and comprehensive evaluations of advanced VLMs in the financial domain. Moreover, fine-tuning Qwen2.5-VL-3B and Qwen2.5-VL-7B on the pyramid-structured question chains enables these models to answer complex financial questions by decomposing them into sub-questions with gradually increasing reasoning demands, yielding average accuracy improvements of 19.52% and 8.06%, respectively, on the dataset.
PyFi: Toward Pyramid-like Financial Image Understanding for VLMs via Adversarial Agents
Yuqun Zhang1∗     Yuxuan Zhao12∗     Sijia Chen1† 1 The Hong Kong University of Science and Technology (Guangzhou) 2 Yantai Research Institute, Harbin Engineering University research@yuqunzhang.com, yx.zhao129@gmail.com, sijiachen@hkust-gz.edu.cn
footnotetext: Equal contribution.footnotetext: Corresponding author.
1 Introduction
Vision language models (VLMs) are capable of step-by-step reasoning, presented as Chain-of-Thought (CoT) Wei et al. (2022), for image understanding in question answering (Q&A) across various scientific domains (Chen et al., 2024a; Yue et al., 2024; Wu et al., 2025d), driven by the development of domain-specific benchmarks (Weng et al., 2025; Wang et al., 2025b; Yue et al., 2024; Wang et al., 2025a), step-wise annotations (Wang et al., 2024c; Zhang et al., 2024b; Qi et al., 2024), and domain-tailored VLMs (Yang et al., 2023; Huang et al., 2024; Lan et al., 2025). In contrast, advancements in the financial domain remain limited. This is because, in finance image understanding, the creation of CoT reasoning datasets, which are highly necessary for the evaluation and fine-tuning of VLMs, requires a high level of expertise and domain knowledge and is therefore scarce.
Figure 1: Overview of the PyFi-600K dataset within our PyFi framework. The dataset is structured as a pyramid comprising 6 capability levels, $1$ financial image types, and 17 financial themes.
To address these issues, previous literature (Chen et al., 2021) collects Q&A samples from financial documents manually, a process that is prohibitively time-consuming and difficult to scale. As a result, subsequent works propose datasets by merging extensive existing financial datasets (Shah et al., 2022; Xie et al., 2024). These repurposed datasets may introduce accumuated biases and inconsistencies inherited from the original sources. More recently, benchmarks such as (Gan et al., 2025; Kim et al., 2024; Luo et al., 2025; Lu et al., 2025; Li et al., 2025a; Liu et al., 2025d) leverage trained VLMs to generate Q&A pairs augmented with CoTs (Xue et al., 2025). However, these samples may contain factual errors or hallucinations, which can distort VLM evaluation and fine-tuning and thereby undermine their reliability in real-world financial applications, where precision and expert-level understanding are essential.
Therefore, this paper proposes a framework named PyFi, which supports comprehensive evaluation and advanced fine-tuning of VLMs through a pyramid-like hierarchy for financial image understanding. Specifically, PyFi includes a synthesized dataset, PyFi-600K shown in Figure 1, comprising 600 thousand (K) samples with step-wise annotations that provide expert-level guidance on how to progressively solve complex financial problems across $17$ categories. The construction of this dataset is enabled by PyFi-adv, a multi-agent adversarial mechanism designed to automatically synthesize and refine samples using expert-level financial knowledge.
PyFi-600K offers three key benefits for the interpretability and fine-tuning of VLMs in finance. First, the samples are organized in a pyramid structure, where the required image understanding capability increases progressively — from basic visual perception at the bottom level to complex financial decision-making at the top. This hierarchy, with capability levels ranging from one to six, enables comprehensive evaluation of VLMs across multiple aspects and supports ranking the capabilities required by VLMs in different financial themes. Second, samples across different levels are interconnected through question chains, demonstrating how questions at higher levels can be addressed step-by-step by leveraging information and reasoning from lower-level Q&A samples. Finally, it supports process supervision (Lightman et al., 2023), as each sample is assigned a reward score that reflects the correctness of its reasoning process in solving the financial problem.
The construction of such a scalable dataset is enabled by our proposed PyFi-adv, a multi-agent adversarial mechanism in which a challenger agent progressively synthesizes increasingly difficult questions, while a solver agent generates answers by reasoning with financial knowledge. The two agents compete within a Monte Carlo Tree Search (MCTS) framework: the challenger begins with simple questions at level 1, either selecting existing ones or generating new ones, and presents them to the solver for response. Based on the answer, the challenger then formulates a more challenging question in the next round. This iterative process continues until the information accumulated suffices to address the decision-making task at level 6.
Our extensive experiments with PyFi-600K reveal that 15 well-known VLMs struggle to answer complex financial questions, with average accuracy dropping from 71.80% at level 1 to 32.95% at level 6. With question chains, we observe that addressing a Level-6 financial decision-making question typically requires correctly answering an average of $10.48$ sub-questions from lower levels. In particular, our results highlight that errors in calculation analysis are a primary cause of VLMs’ failure in final financial decision-making. More importantly, fine-tuning Qwen-VL models on PyFi-600K using question chains improves accuracy by an average of 13.79%, with gains reaching up to 19.52% in the best case.
2 Related Work
Financial Benchmark. The evaluation of Vision-Language Models (VLMs) in the finance domain has seen remarkable progress, from early text-based benchmarks (Chen et al., 2021; Shah et al., 2022; Xie et al., 2024) to more recent efforts targeting financial image understanding (Kim et al., 2024; Zhu et al., 2025a; Shu et al., 2025; Li et al., 2025a; Liu et al., 2025d; Rangapur et al., 2025; Zhao et al., 2025; Gan et al., 2025; Luo et al., 2025; Xue et al., 2025; DENG et al., 2025). However, the question-answering datasets of most benchmarks, such as FinMME (Gan et al., 2025), rely on human annotation, making them costly and difficult to scale. Besides, they fail to support comprehensive assessment of VLMs in the financial domain. However, we introduce an automatically synthesized dataset for the fine-grained, hierarchical evaluation of financial visual understanding across diverse image types and financial themes. More related background can be found in Appendix A.
Financial Large Models. Previous efforts primarily centered on text analysis using financial LMs (Huang et al., 2023; Wu et al., 2023; Yang et al., 2023; Xie et al., 2023; Liu et al., 2025c). Recent financial VLMs (Wang et al., 2023; Bhatia et al., 2024; Huang et al., 2024; Park et al., 2025) integrate visual inputs to facilitate step-by-step reasoning, thereby enabling broader financial decision-making. However, these models lack the expertise and interpretability needed for challenging financial themes, a limitation stemming from the absence of detailed, step-wise problem-solving processes grounded in finance-specific knowledge. In contrast, our dataset contains questions that illustrate how to address financial problems through a progression from simple to complex image understanding and from basic to increasingly deep financial knowledge.
Multi-Agent LM Framework. Recent advances in LMs have enabled Agentic AI to excel in real-world applications, with multiple LM-based agents now collaborating to perceive, learn, reason, and act, supporting intricate tasks such as simulation (Lin et al., 2025; ShengbinYue et al., 2025), reasoning (Liu et al., 2024a; Wu et al., 2025a; Du et al., 2024), trading (Li et al., 2025b; Yu et al., 2024; Wu et al., 2025c), and data synthesis (Tang et al., 2024; Liu et al., 2025a; Xuan et al., 2025). However, to our knowledge, no existing method in finance leverages a multi-agent adversarial approach to synthesize data. In contrast, inspired by Generative Adversarial Networks (GANs) Goodfellow et al. (2014), we present the first method that employs adversarial agents to automatically synthesize hierarchical, fine-grained financial datasets.
3 Framework
Figure 2: Overview of PyFi-adv: a challenger agent competes with a solver agent under the Monte Carlo tree search (MCTS) paradigm to generate question chains that progressively probe deeper capability levels in financial visual reasoning. See the complete chain in Figure 12 in Appendix.
3.1 Financial Dataset in Pyramid Structure
We release PyFi-600K, a large-scale finance dataset comprising 600 thousand (K) samples, each consisting of a question-answer pair, constructed within a pyramid structure that encompasses six levels of finance-specific visual cognition: Perception (PP), Data Extraction (DE), Calculation Analysis (CA), Pattern Recognition (PR), Logical Reasoning (LR), Decision Support (DS). More importantly, PyFi-600K introduces sample chains that demonstrate how to address higher-level pyramid questions by leveraging knowledge from lower-level samples.
Specifically, we denote each finance sample as $\bm{s}^{l}=\left(I,Q,A,r\right)$, where $r\in\left[0,1\right]$ is the reward reflecting its correctness, $I$ is the finance image to understand, $l\in\left[1,6\right]$ indicates the pyramid level to which the question $Q$ and the answer $A$ belong. The sample complexity degree is $C\left(\bm{s}^{l}\right)\in\left[1,5\right]$. And, as presented in right sub-figure of Figure 2, we define sample chain as $\bm{S}^{l}_{1\dots n}=\left[\bm{s}^{1}_{1},...,\bm{s}^{l}_{n}\right]$ where $n$ is the length and $n\geq l$ because samples within the chain can come from the same level. Considering the chain, the sequence of questions increases in levels or complexity degrees, as the answer for each one requires more insightful image understanding and financial expertise than the preceding ones. That is, for any $\bm{s}^{h}_{i}$ and $\bm{s}^{k}_{j}$ with $i,j\in[1,\dots,n]$, $h,k\in[1,\dots,l]$, $i<j$, we have that $h\leq k$, and if $h=k$, then $C\left(\bm{s}^{h}_{i}\right)\leq C\left(\bm{s}^{k}_{j}\right)$. Subsequently, the reasoning process for answering the question in any sample within the chain depends on its predecessor samples. Specifically, when $\bm{s}^{l}_{j}$ is addressed by a VLM parameterized by $\bm{\theta}$, we have $p_{\bm{\theta}}\left(A|I,Q\right)<p_{\bm{\theta}}\left(A|I,Q,\left\{\bm{s}^{l}_{i}\right\}^{j}_{i=1}\right)$, meaning that the reasoning information and financial knowledge in answers from predecessor samples enable more reliable question answering for the current sample.
Therefore, compared to existing manual and automatic evaluation dataset construction methods, we observe the following three advantages of the proposed approach for the finance domain:
Hierarchical Cognition in the dataset embodies human-level problem solving processes, progressing from basic perception to complex financial decision-making. Samples at each level $l$ correspond to a specific cognitive capability, which builds upon those from lower levels i=1..,l1i=1..,l-1 or complexities. Thus, evaluating VLMs on these hierarchically structured samples enables a comprehensive assessment of their abilities in the finance domain. That is, any financial reasoning performed for questions from higher levels, which require rare, expert-level knowledge, relies on foundational financial questions involving base cognitions such as CA. Besides, we can identify how VLMs perform, whether better or worse, at each cognition level during financial reasoning, thereby revealing which level is crucial for which financial questions and to what extent.
Progressive Logical Chain in the dataset is presented as the sample chain $\bm{S}^{l}_{1\dots n}$ shown in Figure 2, which demonstrates how a challenging financial question at level $l=6$ is solved by progressively addressing a sequence of simple-to-complex sub-questions. This aligns with financial decision-making logic, where each step builds upon the previous one to form a cumulative and logically coherent reasoning flow. Such a traceable chain enhances the interpretability and justifiability required in the finance domain. Moreover, VLMs fine-tuned on such chains are capable of solving complex financial problems by accumulating knowledge through first addressing a series of simpler problems.
Step-wise Reward Score $r$ of each sample reflects its reliability; thus, the scores along a sample chain support process supervision (Lightman et al., 2023) in reasoning with financial image understanding. Using these scores, we enable the training of process reward models that act as verifiers for any financial VLM.
3.2 Automatic Synthesization with Adversarial Agents
Existing financial datasets rely solely on human annotation, while automated methods from other domains, such as rStar-Math (Guan et al., 2025) for mathematical reasoning, fail to meet the requirements of finance, where interpretability and reliable reasoning are essential, along with increasing complexity, domain expertise, and hierarchical cognition. We achieve this by proposing a multi-agent adversarial mechanism, referred to as PyFi-adv, which constructs question chains, as illustrated in Figure 2.
Specifically, PyFi-adv contains a challenger agent that is supported by an off-and-shelf VLM, denoted as $\psi$, and similarly, a solver agent $\phi$. Inspired by the adversarial mechanism in AlphaGo (Silver et al., 2016), our two agents compete within an MCTS framework to automatically synthesize high-quality reasoning samples. In the finance domain, this approach substantially outperforms existing automatic data augmentation methods discussed in subsection A.3.
Confrontation for Sample Synthesization, as shown by the Gray oval circle in Figure 2, is that when $\phi$ aims to correctly answer the given question, $\psi$ checks the answer to propose a more challenging follow-up question to hinder the solver. This is formulated as $Q_{i}\sim p_{\psi}\left(Q_{i}|I,{\bm{\mathsfit{P}}}_{q},\bm{S}^{l}_{1\dots i-1}\right)$ and thus $A_{i}\sim p_{\phi}(A_{i}|I,Q_{i},{\bm{\mathsfit{P}}}_{a},\bm{S}^{l}_{1\dots i-1})$, where ${\bm{\mathsfit{P}}}_{q}$ and ${\bm{\mathsfit{P}}}_{a}$ are the agent prompts.
Exploration for Diversity Guarantee aims to prevent the adversarial process from converging to a set of questions with identical content. At each step, a Bernoulli sampling is used to determine the $\psi$ whether to explore a new question path or exploit an existing one. For exploitation, the UCT formula is employed to select the most promising prior question.
Expansion for Hierarchical Questions is naturally achieved through the adversarial agents: once $\psi$ performs a question expansion, it generates a subsequent, more complex question based on the accumulated $\bm{S}^{l}_{1\dots i}$. Agent $\phi$ then produces an answer $A_{i+1}$, forming a new sample $\bm{s}^{l}_{i+1}$. Thus, a hierarchical question structure is built through repeated adversarial interactions that continue until either the top-level ($l=6$) question is addressed or a predefined depth limit is reached.
Backpropagation for Reward Assignment adjusts the success rate of each sample by backpropagating the comparison between the answer of the top-level question and the ground truth down to individual samples in the chain, as indicated by the orange dotted line in Figure 2. Upon convergence of the MCTS, the resulting success rate is used as the reward score, providing a correctness label that measures the reliability of the synthesized dataset.
4 Experiments
4.1 Setup
Table 1: Evaluation of 15 Pre-trained VLMs on PyFi-600K across 6 capability levels and 5 complexity degrees, conducted on the Evaluation Dataset after Exclusion. This dataset is derived from an initial random selection of 1000 instances from PyFi-600K, followed by an additional filtering step to address potential data leakage. The final test set consists of 301 samples. For simplicity, we only present the results from complexity degrees 3, 4, and 5. The full model names listed in the "Model" column are: GPT-4.1, InternVL3-38B, Claude-opus-4-1-20250805, DeepSeek-VL2, Hunyuan-Large-Vision, ERNIE-4.5-turbo-vl, Moonshot-V1-128k-Vision-Preview, Moonshot-V1-32k-Vision-Preview, Moonshot-V1-8k-Vision-Preview, Qwen3-VL-Plus, Qwen2.5-VL-72B-Instruct, Qwen2.5-VL-32B-Instruct, Qwen2.5-VL-7B-Instruct, and Qwen2.5-VL-3B-Instruct.
Model Overall PP DE CA PR LR DS 33 44 55 GPT-4.1 (OpenAI, 2025) 52.99 86.84 47.64 40.62 57.14 54.84 38.46 52.63 46.55 53.85 InternVL3 (Zhu et al., 2025b) 52.91 76.32 56.37 44.53 43.88 53.23 30.77 41.78 44.83 67.31 Claude-opus (Anthropic, 2024) 64.70 80.92 62.74 66.02 58.16 72.58 32.69 66.12 55.60 73.08 GLM-4.5V (Hong et al., 2025) 74.75 89.47 78.07 65.23 75.51 75.81 46.15 72.37 64.22 73.08 DS-VL2 (Wu et al., 2024b) 45.18 76.32 44.58 36.33 42.35 42.74 19.23 34.54 42.24 46.15 Hunyuan-LV (Sun et al., 2024) 59.72 90.79 58.96 36.72 63.27 69.35 51.92 50.00 57.76 80.77 ER-4.5-TVL (Baidu-ERNIE-Team, 2025) 34.47 34.87 33.96 36.72 25.00 45.16 36.54 38.82 25.43 65.38 MV1-128k (Team et al., 2025b) 54.57 80.92 61.08 39.06 41.84 58.87 38.46 46.71 43.97 57.69 MV1-32k (Team et al., 2025b) 54.40 80.26 61.56 37.11 40.31 62.90 38.46 45.72 43.97 61.54 MV1-8k (Team et al., 2025b) 54.90 80.26 61.79 39.84 42.35 58.06 38.46 48.03 43.97 55.77 Q3-VL-Plus (Yang et al., 2025a) 51.00 77.63 54.25 37.50 41.84 57.26 32.69 45.72 37.50 65.38 Q2.5-VL-72B (Bai et al., 2025) 48.84 70.39 59.91 33.20 35.71 46.77 26.92 36.51 31.03 57.69 Q2.5-VL-32B (Bai et al., 2025) 43.19 60.53 45.28 37.50 36.73 42.74 28.85 39.14 30.60 59.62 Q2.5-VL-7B (Bai et al., 2025) 37.87 54.61 49.53 34.77 18.37 24.19 15.38 31.91 17.67 38.46 Q2.5-VL-3B (Bai et al., 2025) 20.51 36.84 15.57 21.48 16.84 21.77 19.23 23.03 18.97 30.77
Datasets. PyFi-600K dataset consists 603,859 samples and 62,660 sample chains with reward scores of 1. For evaluation, we proportionally sample 1,000 samples from each capability level to construct the test set across all cases. See detail in Appendix B.
Data Leakage Mitigation. To ensure a fair evaluation, we systematically identify and exclude samples that exhibit data leakage. Specifically, any question that can be correctly answered by more than half of the evaluated models without the image is considered leaked. This reduces our test set from 1,000 to 301 samples. Detailed analysis and justification are provided in Appendix D.
Models. For evaluation, we use 15 VLMs as listed in Table 1. For SFT, we fine-tune Qwen2.5-VL-3B-Instruct and Qwen2.5-VL-7B-Instruct on sample chains to obtain PyFi-QwenVL-3B and PyFi-QwenVL-7B, respectively.
SFT. We fine-tuned the models on approximately 47K sample chains using two dataset variants: one with chain-of-thought (CoT) annotations (w/ CoT) and one without (w/o CoT). In the CoT variant, each question chain is converted into a CoT reasoning sequence, where each reasoning step corresponds to a sub-question and its answer. Thus, a chain with $n$ samples yields a CoT sequence with $n$ reasoning steps. In the non-CoT variant, only the question and answer from the final sample in the chain are used. We use the AdamW optimizer with a learning rate of $1.0\times 10^{-4}$, cosine learning rate scheduling, and a warmup ratio of 0.1. Training is conducted for one epoch with an effective batch size of 8. For parameter-efficient fine-tuning, we apply LoRA with full-module adaptation. All experiments are carried out on four NVIDIA RTX 5090 GPUs.
[…]
4.2 Main Results
Evaluations of $15$ Pre-trained VLMs are presented in Table 1. In summary, we observe that as the capability level and complexity of financial image understanding increase, the accuracy of VLMs gradually declines. Specifically, the average accuracy across all models decreases from 71.80% at level 1 (Perception) to 32.95% at level 6 (Decision Support). This indicates that while current VLMs can handle basic visual perception in the financial image understanding, they struggle with complex financial questions that require higher-level cognitive abilities. Notably, GLM-4.5V achieves the highest overall accuracy of 74.75%, demonstrating its strong performance across various capability levels. However, even this model’s accuracy drops to 46.15% at level 6, highlighting the challenges in financial decision-making. Most models, including GPT-4.1 and InternVL3-38B, achieve over 70% accuracy at level 1 but fall below 40% at level 6, resulting in a relatively low average overall accuracy of 50%.
Evaluations of Fine-tuned Qwen-VL models are presented in Figure 3. In summary, with question chains arranged in a pyramid manner, the models, especially the smaller ones, show a significant increase in accuracy and gain the ability to reason progressively, moving from low-level to high-level questions until they reach the solution, as shown in Figure 5. Specifically, both our PyFi- models exhibit substantial accuracy gains. PyFi-CoT related models improves the 3B model by 19.52% points and the 7B model by 8.06% points over their local baselines. The gain is notably larger for the smaller 3B model. Notably, on higher-level financial questions, such as Decision Support (DS), PyFi-CoT ones improve the 3B model by 9.62% and the 7B model by 23.08%. The substantially larger gain for the 7B VLM suggests that larger VLMs benefit more from question-chain fine-tuning in challenging financial scenarios. Notably, in Pattern Recognition (PR) and Logical Reasoning (LR), PyFi-QwenVL-3B even surpasses the accuracy of the much larger 72B VLM, as shown in Table 1. Building on the detailed comparison in Figure 5, we argue that sample-chain fine-tuning equips VLMs with the ability to solve complex financial problems by iteratively addressing a sequence of self-generated sub-questions of increasing capability levels, thereby significantly improving interpretability and reliability.
Figure 3: Comparison between Qwen2.5-VL models and ours ( PyFi-QwenVL-3B and PyFi-QwenVL-7B).
Figure 4: Evaluation of error proportions across capability levels leading to incorrect financial decisions. As illustrated in the right sub-figure, the sample chain enables error tracing: when Claude-Opus-4-1 fails to answer the Level 6 financial decision-making question, we can trace back to identify the specific lower-level sub-questions that were answered incorrectly, thereby pinpointing the exact level at which the error originated.
4.3 Main Benefits and Insights
Evaluations in Table 1 and Figure 3 enable a fine-grained analysis of how — and to what extent — VLMs meet the varying capability requirements of financial image understanding across different levels of complexity. The phenomenon of decreasing accuracy as the level increases in Table 1 indicates that existing VLMs can only meet the most basic image understanding in finance and are difficult to be used for complex financial decision-making. For instance, while models like GLM-4.5V perform well on basic perception questions (level 1), their accuracy drops considerably on decision support questions (level 6). This trend is consistent across most models, suggesting that greater efforts are needed to enhance VLMs’ capabilities in the finance domain. Thus, the results in Figure 3 demonstrate the importance of our question-chain fine-tuning: Qwen models learn to make accurate financial decisions by first addressing a series of sub-questions in a progression from simple to complex, a process aligned with human cognition.
Tracing back from Decision Support to Lower-level Q&As, as illustrated in Figure 4, significantly improves the interpretability of VLMs in financial decision-making. In Figure 4, we first collect Level-6 samples that VLMs fail to answer correctly and analyze their question chains by evaluating VLMs on the questions at each lower level in the chain. In the finance domain, addressing a Level-6 financial decision-making question typically requires correctly answering sub-questions from lower levels. Thus, when an error occurs, as indicated by the red cross in the right sub-figure, PyFi-600K enables us to trace the problem-solving chain from DS to LR and eventually to PP, thereby diagnosing how an error in the DE leads to the final mistake. Such a tracing-back approach is crucial for understanding the limitations of VLMs in financial decision-making, as it reveals at which capability level a failure occurs.
Identifying the Most Frequent Failure Capability Levels, as presented in the left part of Figure 4, reveals that the key to improving VLMs’ performance in the financial domain lies in strengthening their capabilities at the CA (Calculation Analysis) level. Specifically, the pyramid-like structure and the question chain of PyFi-600K facilitates detailed error analysis, allowing us to identify specific levels where VLMs struggle. That is, most VLMs, even the advanced Claude-opus-4-1-20250805 and Qwen3-VL-Plus, fails to answer more than 40% of the question from CA, revealing that the calculation analysis is a critical bottleneck for them. More importantly, by comparing the error proportions between six capability levels, we can prioritize which capabilities to enhance first. For instance, since CA has the highest error proportion, focusing on improving calculation analysis skills in VLMs could lead to the most significant overall performance gains in financial decision-making tasks. In addition, this comparison also shows that, in the finance domain, analyzing numerical calculations from images is one of the most critical capabilities for accurate decision-making, even more important than logical reasoning, aligning with conclusions drawn by real-world finance experts.
Figure 5: Qualitative analysis of level-wise COT by comparing the Qwen2.5-VL-3B-Instruct and PyFi-QwenVL-3B in the challenging financial decision-making.
Addressing Financial Decision-Making with Level-wise CoT, as depicted in Figure 5, is not only reliable but also highly interpretable. Specifically, we compare the responses of Qwen2.5-VL-3B-Instruct and PyFi-QwenVL-3B — the latter fine-tuned on question chains— when addressing the financial problem. The Qwen2.5-VL-3B-Instruct produces a short and disorganized reasoning process, reflecting its limited financial reasoning capability. In contrast, PyFi-QwenVL-3B follows a clear, structured, and logical reasoning flow, which is essential for expert-level performance in finance. For instance, our model first addresses basic PP questions, then proceeds to DE and CA, and finally tackles high-level LR and DS. This level-wise CoT not only improves answer reliability but also provides a transparent and traceable reasoning path that users can easily follow and verify, making it especially valuable in financial contexts where stakes are high. Moreover, experts can inspect this professional decision-making process and provide detailed and direct feedback to further refine the model.
Providing Process Supervision, in the form of a reward score for each sample in our PyFi-600K, enables research into training reliable verifier models, known as process reward models (PRMs), to facilitate checking the correctness of each step generated by financial VLMs. Even though, due to space limitations, this paper does not perform such training, we use the reward scores to filter out low-quality samples during both evaluation and fine-tuning.
Mitigating Hallucinations. As shown in Figure 5 and Figure 4, our support of Level-wise CoT in training effectively mitigates model hallucination by enforcing a grounded reasoning process: high-level conclusions (e.g., DS) are derived strictly from verified low-level findings (e.g., PP and DE), rather than from ungrounded speculation. This structure ensures traceability and verifiability, as every reasoning step can be explicitly traced back to concrete evidence in preceding steps.
5 Conclusion
In this paper, we have proposed a framework named PyFi, containing the dataset PyFi-600K and the multi-agent adversarial mechanism PyFi-adv, to facilitate pyramid-like financial image understanding for VLMs. With the Q&A samples organized in a pyramid structure, PyFi-600K enables comprehensive evaluation of VLMs across $6$ capability levels, $17$ financial themes and $11$ image types. Our novel level-wise question chain supports the interpretability of the performance of VLMs in financial decision-making by 1) showing importance ranking of capability level in accurate decisions and 2) enabling a back-track of which errors in lower levels led to the decision failure. More importantly, with PyFi-adv, the dataset is scalable without human annotations, meaning the size and quality can be improved continuously. After performing extensive experiments on $15$ VLMs and especially fine-tuning on Qwen models with question chains, we have verified these benefits of the PyFi by showing detailed, hierarchical, and fine-grained results, exposing important insights such as Calculation Analysis (CA) is the most crucial for correct decision-making and fine-tuned small VLMs gain $19.52\%$ accuracy improvement without compromising the model’s existing capabilities on general tasks.
Appendix Contents
Appendix A Related Work Discussion
A.1 VLM Reasoning
Vision-language models (VLMs) have achieved impressive performance on complex tasks by leveraging step-by-step reasoning, decomposing intricate problems into simpler subtasks and solving them sequentially to reach a final solution (Guo et al., 2025; Team et al., 2025b; Bai et al., 2025; Hong et al., 2025; Team et al., 2025a; Wang et al., 2025d). Benchmarking reasoning ability is essential for guiding future advancements. In contrast to earlier methods that evaluated reasoning holistically via a single scalar score, recent work has shifted toward fine-grained assessment, enabling step-by-step validation of the chain-of-thought (CoT) process (Wu et al., 2025d; Chaoyou et al., 2023; Zhu et al., 2024; Chen et al., 2024a; Liu et al., 2024b; Yue et al., 2024; Chen and Song, 2025b). This paradigm has seen significant traction in domains like mathematical reasoning (Wang et al., 2024b; Lu et al., 2024b; Zhang et al., 2024a; Weng et al., 2025; Li et al., 2025c; He et al., 2025; Wang et al., 2025b; Qiao et al., 2024) and code generation (Xiao et al., 2025, 2024; Wang et al., 2025a; Awal et al., 2025; Li et al., 2024; Wu et al., 2024a), where outputs are inherently verifiable and intermediate steps can be automatically annotated.
The success in these domains stems from the inherent verifiability of their outputs, which enables automatic construction of process-annotated datasets for fine-tuning large language models (LLMs). Two types of verifiers are commonly employed: the Outcome Reward Model (ORM), which assigns a scalar reward to the entire solution, and the Process Reward Model (PRM) (Ma et al., 2023; Zhang et al., 2025b; Wang et al., 2025c; Choudhury, 2025; Lambert et al., 2024), which evaluates correctness at each step of the reasoning path. PRM’s step-wise scoring enables precise error localization, facilitating more targeted feedback.
However, this progress has yet to be effectively transferred to vertical domains such as finance. Reasoning in finance typically involves multistep interpretation of charts, tables, and contextual narratives, requiring not only accurate understanding of financial visuals and deep domain knowledge, but also often yielding open-ended answers that are inherently difficult to verify precisely. This ambiguity impedes the generation of dense, step-wise process rewards essential for PRM training.
A.2 Financial Benchmarks
Existing financial benchmarks remain limited: they reduce complex reasoning to single scalar scores, obscure intermediate reasoning steps, and lack support for hierarchical or diagnostic evaluation (Kim et al., 2024; Zhu et al., 2025a; Shu et al., 2025; Li et al., 2025a; Liu et al., 2025d; Rangapur et al., 2025; Zhao et al., 2025; Gan et al., 2025; Luo et al., 2025; Xue et al., 2025; DENG et al., 2025). Additionally, current evaluation datasets consist only of isolated question-answer pairs, making it difficult to uncover hidden relationships within the data (Yang et al., 2025b). Compounding this challenge, these benchmarks rely on costly manual annotations that hinder scalability, and no large-scale synthetic dataset currently exists to train finance-specific preference reward models (PRMs) for visual reasoning tasks.
Specifically regarding scalability, FAMMA-LivePro (Xue et al., 2025), a manually annotated benchmark, contains 103 questions crafted by senior finance experts. These questions are of high quality and accuracy, and capture the complexity of real-world financial scenarios. However, such manual annotation demands highly qualified experts and is both time-consuming and labor-intensive, making it difficult to scale. While some benchmarks (Wang et al., 2024d; Gan et al., 2025; Luo et al., 2025; Lu et al., 2025; Li et al., 2025a; Zhu et al., 2025a; Zhao et al., 2025; Shu et al., 2025; Liu et al., 2024b) explore using large models for automatic data synthesis, they typically prompt these models to directly generate expert-level, complex financial questions. However, due to the difficulty large models face in understanding complex financial visuals, they often rely heavily on provided textual cues, leading to question homogenization and hallucinations. This results in question-answer pairs of inconsistent quality that still require substantial manual verification. On the other hand, most financial benchmarks only release their datasets without disclosing their data augmentation methods, making direct comparison of such methods difficult.
As a result, the finance domain still lacks benchmark datasets capable of supporting hierarchical capability assessment, evaluating step-by-step reasoning, enabling fine-grained reward modeling, and achieving fully scalable data augmentation.
A.3 Automatic Data Augmentation
Existing automated annotation methods fall into two categories. The first includes Math-Shepherd (Wang et al., 2024c), AutoPSV (Lu et al., 2024a), AutoPRM (Chen et al., 2024b), and AdaptiveStep (Liu et al., 2025b), which score or decompose steps of a fixed problem using verifiers or confidence heuristics. While efficient, they are passive and cannot generate new problems or structure them by difficulty. The second category, including OmegaPRM (Luo et al., ), rStar-Math (Guan et al., 2025), GroundedPRM (Zhang et al., 2025a), and RMultiplex200K (Chen and Song, 2025a), leverages MCTS or structured planning to create higher-quality, verifiable paths. Specifically, OmegaPRM employs a divide-and-conquer MCTS algorithm to systematically locate the first incorrect step in a reasoning chain, enabling efficient and high-quality data generation. GroundedPRM also uses MCTS but integrates external validation tools to ensure the fidelity of each step, resulting in a more data-efficient PRM that achieves SOTA performance with far fewer annotations.
While existing methods such as Math-Shepherd, AutoPSV, AutoPRM, and AdaptiveStep effectively annotate steps of a given problem through passive, single-agent verification, they are inherently limited to flat, post-hoc labeling and cannot generate new problems or structure them by difficulty. Even MCTS-based approaches though capable of constructing high-fidelity reasoning paths—focus exclusively on verifying or refining a single solution trajectory. In contrast, our work introduces an active, adversarial two-agent framework embedded within the MCTS architecture: one agent (the questioner) generates financial reasoning problems, while the other (the answerer) solves them, with the questioner progressively increasing difficulty based on the answerer’s performance.
This co-evolutionary process is specifically designed for the financial domain, where reasoning often requires layered knowledge, from basic accounting principles to complex risk modeling, and where errors must be interpretable and traceable. Our method explicitly constructs a hierarchical, interpretable curriculum of financial problems, in which each subsequent question builds upon prerequisite concepts and demands deeper domain understanding. Non-MCTS methods lack the structured search mechanism needed to jointly optimize problem generation, difficulty progression, and financial knowledge dependency. MCTS, by contrast, provides the necessary planning backbone to balance exploration of novel financial scenarios with exploitation of solvable, pedagogically meaningful chains. Thus, MCTS is essential not only for scalability but also for realizing a generative, knowledge-aware annotation process that produces explainable, difficulty-graded financial reasoning data, extending MCTS’s role beyond verification to structured curriculum synthesis tailored to domain-specific reasoning.
A.4 Our Contributions
To address above shortcomings, we aim to: (1) comprehensively evaluate VLMs across a hierarchy of cognitive levels, from basic perception to expert financial decision-making, enabling granular diagnosis of strengths and weaknesses at each reasoning stage and revealing which foundational cognitions are critical for higher-level financial tasks, and uncover meaningful data relationships; (2) provide traceable, progressive logical chains that mirror real-world financial reasoning, enhancing interpretability, explainability, and justifiability while enabling VLMs to learn complex problem-solving through cumulative, step-by-step knowledge building; (3) introduce step-wise reward scores for each reasoning step, facilitating automated, scalable verification via Process Reward Models and overcoming the limitations of costly manual annotation, and (4) propose an automatic data augmentation method that decomposes the task of generating complex financial question-answer pairs into incremental process, starting from simple visual perception questions and progressively increasing difficulty and financial reasoning depth, all without requiring human verification. Finally, we present a comparison between our PyFi-600K dataset and existing evaluation datasets in Table 2.
Table 2: Comparison of PyFi-600K with Existing Evaluation Dataset. In Multimodal column, "T" means Text modality, "I" means Image modality. In the Volume column, K is used as the unit.
Dataset Volume Multimodal Automatic Synthesized Hierarchical Cognition CoT Step-wise Reward General Domain MME (Chaoyou et al., 2023) 2.4 T & I MMDocBench (Zhu et al., 2024) 4.3 T & I MMStar (Chen et al., 2024a) 1.5 T & I MMBench (Liu et al., 2024b) 3.2 T & I MMMU (Yue et al., 2024) 11.6 T & I Finance Domain FinQA (Chen et al., 2021) 8.3 T FLUE (Xie et al., 2023) 26.3 T FLARE (Xie et al., 2023) 19.2 T CF-Benchmark (Lei et al., 2023) 3.9 T BizFinBench (Lu et al., 2025) 7.0 T FinMME (Luo et al., 2025) 11.1 T & I MME-Finance (Gan et al., 2025) 1.2 T & I FAMMA-Basic (Xue et al., 2025) 1.9 T & I CFBenchmark-MM (Li et al., 2025a) 9.7 T & I FinMMR (Tang et al., 2025) 4.3 T & I MultiFinBen-Vision (Peng et al., 2025) 21.8 T & I PyFi-600K (Ours) 600 T & I
Appendix B Data Details
B.1 Data Statistics
The complete PyFi-600K dataset comprises 603,859 samples, each comprising a financial image paired with a corresponding question-answer (Q&A) pair, along with essential background information (caption, theme, acronym expansions) of the image and a summary analysis of its context. Each sample is annotated with a capability level and complexity rating. Specifically, Level 1 (Perception) contains 73,006 samples (12.09% of the total); Level 2 (Data Extraction) includes 159,601 samples (26.43%); Level 3 (Calculation Analysis) comprises 90,229 samples (14.94%); Level 4 (Pattern Recognition) has 109,848 samples (18.19%); Level 5 (Logical Reasoning) accounts for 117,751 samples (19.50%); and Level 6 (Decision Support) includes 53,424 samples (8.85%). Notably, Level 2, Data Extraction, is the most prevalent, underscoring its foundational role in complex reasoning tasks. The dataset also includes over 60,000 sample chains, whose lengths vary from 1 to 26 samples, with an average chain length of 11.48. Question lengths range from 29 to 853 characters, with an average length of 142 characters, as shown in Table 3. The full distribution across capability levels and complexity degrees is shown in Table 4.
Table 3: Question Length and Question Chain Length Statistics
Statistic Question Length (Characters) Chain Length Mean 142.11 11.48 Median 137.00 12 Minimum 29 1 Maximum 853 26 Std Dev 50.54 2.55
Table 4: Distribution of Question Capability Levels and Complexity Degree
Classification Type Count Percentage (%) Capability Level 1 Perception 73,006 12.09 2 Data Extraction 159,601 26.43 3 Calculation Analysis 90,229 14.94 4 Pattern Recognition 109,848 18.19 5 Logical Reasoning 117,751 19.50 6 Decision Support 53,424 8.85 Complexity Degree 1 91,143 15.09 2 148,465 24.59 3 165,072 27.34 4 139,654 23.13 5 59,527 9.86 Total 603,859 100.00
The distribution of question chain lengths by segment has shown in Table 5. The results indicate that the majority of sample chains (84.47%) fall within the 11-15 length range, suggesting that most tasks in the dataset require a moderate to substantial number of reasoning steps. Since chain length reflects the number of logically sequential question-answer pairs needed to arrive at a final answer, longer chains correspond to higher task complexity. The scarcity of very short (1-10, 15.11%) or very long (16-26, 0.42% combined) chains further implies that the dataset is intentionally curated to focus on multistep reasoning scenarios that are neither trivial nor excessively extended.
Table 5: Distribution of question chain lengths by segment
Length Range Count Percentage (%)
1–10 9,456 15.11
11–15 52,863 84.47
16–20 241 0.39
21–26 21 0.03
We also observed that, generally, the average question length increases with higher capability levels, indirectly reflecting a progressive rise in task difficulty, as shown in Figure 6. However, an unexpected decrease occurs at Level 4 (Pattern Recognition). This may be because pattern recognition tasks often require identifying and interpreting intricate patterns rather than relying on simple textual matching, which could lead to more concise yet semantically dense questions that challenge model performance.
Figure 6: Capability Level with Question Length.
Finally, we analyzed word frequency in the questions, as shown in Table 6 and Figure 7. Excluding common function words such as prepositions and basic terms like "and" or "where," the most frequently occurring words include "price", "growth", "trend", "GDP", "debt", "investment", "income", "share", "revenue", "change", "risk", "government", and "market". This vocabulary highlights the dataset’s rich thematic coverage across diverse domains of finance.
Table 6: Top most frequent words in the questions
Rank Word Frequency Percentage (%) 1 price 112,502 1.31 2 growth 99,789 1.17 3 trend 83,465 0.97 4 gdp 63,960 0.75 5 debt 42,057 0.49 6 investment 38,931 0.45 7 income 34,923 0.41 8 share 33,044 0.39 9 revenue 29,665 0.35 10 change 29,019 0.34 11 risk 28,486 0.33 12 government 27,847 0.33 13 market 26,763 0.31
Figure 7: Top most frequent words in the questions.
Before fine-tuning, we analyzed the answer distribution across all samples in the question chains and observed significant category imbalance. As shown in Table 7, answer option A accounts for 43.39% of all responses, while B makes up 34.20%. In contrast, options C and D represent only 15.81% and 6.08%, respectively, with options E-J collectively comprising less than 0.52%. This skew suggests a generation bias, likely introduced during synthetic question creation, where the model tends to place the correct answer in the first two choices. To mitigate this bias and ensure a fair evaluation, we manually rebalanced the answer distribution in a subset of 1,000 test samples so that options A, B, C, and D each appear with approximately equal frequency. During fine-tuning, we then compared model accuracy under the original imbalanced distribution versus the rebalanced version to quantify the impact of answer-position bias on model performance. This controlled comparison helps isolate whether apparent gains in accuracy stem from genuine reasoning ability or merely from exploiting skewed answer patterns.
Table 7: Answer distribution in the Question Chain with non-standard responses grouped as "Wrong Format"
Answer Category Percentage (%) A 43.39 B 34.20 C 15.81 D 6.08 E 0.44 F 0.06 G <<0.01 H <<0.01 I <<0.01 J <<0.01 Wrong Format <<0.01 Total 100.00
B.2 Instructions
Data Privacy
Data is collected from open-source and publicly available resources, including reports, textbooks, and online documents. When using MinerU (Wang et al., 2024a) for data extraction, we strictly adhere to the copyright and licensing terms of each source. Furthermore, we provide comprehensive citation information for every financial document in our collection, including details such as title, author, and publisher. Most importantly, the dataset is used exclusively for research purposes and involves no financial gain.
Data Format
All textual data in PyFi-600K are formatted in JSON, while images are stored in .jpg format. We verify that each sample contains exactly one image, belongs to a specific level within the pyramid hierarchy, and is part of one or more sample chains.
Data Structure
The complete dataset is stored in a tree structure represented in JSON format. The root node of the tree is a placeholder with no semantic meaning, while every other node represents a Q&A pair. Each branch extending from the root to a leaf node constitutes a complete sample chain. Nodes are assigned to different levels of a capability pyramid based on their hierarchical competence. Each node contains the following fields: the question, the answer, an associated image, essential visual context required to answer the question, and a summary/analysis of the contextual information.
B.3 Data Processing
To construct the final PyFi-600K dataset, we adopt a concise multi-stage pipeline: collect public financial documents; parse PDFs to extract images and associated text; evaluate compliance and classify images (normal/abnormal/extremely abnormal), enhancing only normal ones with contextual metadata; categorize images into content themes and chart types; maintain provenance; and sample by relevance and complexity with expert validation. An overview is shown in Figure 8.
Figure 8: Overview of Data Processing Pipeline.
We aggregated over 30,000 public financial PDFs in Chinese and English via web crawling and institutional APIs, then applied technical checks (corruption, password protection, readability) and domain relevance filters with curated inclusion/exclusion keywords to retain about 10,000 documents. We leverage MinerU (Wang et al., 2024a) to parse validated PDFs and extract images, text blocks, and layout metadata, preserving image–caption associations with page indices and bounding coordinates. Extracted assets were reorganized into a standardized hierarchical directory by document and page (separating images and contextual text) for efficient batch processing and traceability. We screened approximately 820,000 images using a VLM-guided dual 0–10 scoring for compliance (financial visualization criteria) and complexity (visual/data density and structural sophistication), generated per-image JSON summaries, and conducted a targeted manual audit. Images were labeled as normal/abnormal/extremely abnormal by combining caption-match windows around image tags with nearby-image proximity signals; only normal images were enhanced with contextual metadata. We follow the following criteria to select the target images for our study:
  • Financial trend data chart only; no non-informative elements.
  • Clear, legible Chinese or English text; unambiguous recognition.
  • No obvious errors or omissions; follows financial data conventions.
  • Clearly highlights dynamic financial processes.
  • Tables/grids are secondary and must not dominate.
  • Explicit relationships shown via visual connections.
  • Addresses at least one finance-related issue.
  • Universally discernible; self-contained composition.
Context Extraction. The retained images are classified into three status categories (normal, abnormal and extremely abnormal) based on their contextual clarity within the document. A normal image is defined by a clear one-to-one match with a single caption located nearby and no other adjacent images, indicating reliable and unambiguous correspondence. Only images classified as normal are enriched with detailed contextual metadata from the surrounding text, providing each with explanatory insights.
An abnormal image arises when there is contextual ambiguity, either because the image has other images close to it or because multiple captions appear near it. An extremely abnormal image is identified when no caption is found within the search window, signaling a high risk of missing context. This classification ensures that contextual enhancement is applied only to images with a well-defined and correct caption relationship.
Topic Classification. We utilize the VLM to classify each image into one of 17 distinct financial content themes and one of 11 chart types, based on both the visual content of the image and its extracted contextual information.
Content Theme and Chart Type Classification
I. Content Theme
This dimension describes the core financial subject of the image, including Macroeconomic Indicators, Financial Markets & Products, Commodities & Real Estate Markets, Bonds & Fixed Income, Monetary & Fiscal Policy, International Trade & Capital Flows, Corporate Finance & Valuation, Industry Analysis, Investment Theory & Portfolio Management, Risk Models & Management, Economic Cycles & Market Theories, Microeconomic Principles, Demographics & Socioeconomics, Financial Systems & Infrastructure, Geospatial Economic Data, Organization & Regulation, and Financial History & Documentation, as shown in Figure 9 and Figure 10.
II. Chart Type
The chart types including Line Chart, Bar Chart & Column Chart, Pie Chart & Donut Chart, Scatter Plot & Bubble Chart, Table, Diagram & Schematic, Radar Chart, Heatmap, Candlestick Chart, Photograph, and Infographic, as shown in Figure 11.
Macroeconomic Indicators
Financial Markets & Products
Commodities & Real Estate Markets
Industry Analysis
Monetary & Fiscal Policy
International Trade & Capital Flows
Corporate Finance & Valuation
Bonds & Fixed Income
Investment Theory & Portfolio Management
Risk Models & Management
Figure 9: Overview of content themes.
Economic Cycles & Market Theories
Geospatial Economic Data
Demographics & Socioeconomics
Financial Systems & Infrastructure
Microeconomic Principles
Organization & Regulation
Financial History & Documentation
Figure 10: Overview of content themes (continued).
Line Chart
Pie Chart & Donut Chart
Table
Bar Chart & Column Chart
Scatter Plot & Bubble Chart
Candlestick Chart
Diagram
Figure 11: Overview of chart types. In images used for generating question-answer pairs, Radar Charts, Heatmaps, and Photographs are rarely used.
Reference Info Extraction. We employ a LLM to extract bibliographic references and citation metadata from financial and academic documents. The model parses footnote, endnotes, and reference sections to identify cited works and extract key fields, including author(s), publication venue, year, and DOI (when available), while mapping each citation to its contextual use within the text.
Data Sampling. The image sampling strategy is structured around four key dimensions: complexity degree, compliance level, content theme, and chart type. The compliance level (ranging from 0 for non-compliant to 10 for fully compliant) reflects the degree to which the image satisfies the specified financial data visualization criteria. The complexity degree (from 0 for very simple to 10 for very complex) is assigned based on visual intricacy, data density, and structural sophistication. Each image is tagged with at least one of 17 content themes and at least one of 11 image types.
From an initial pool of 820,816 images, we applied a two-stage filtering process to construct a comprehensive financial evaluation dataset. First, we retained only images with compliance levels 9 and 10, indicating full or near-full adherence to the specified financial data visualization criteria. Non-financial images were excluded.
From the compliant subset, we selected the top 20,000 images with the highest complexity degree, reflecting visual intricacy and informational density. We further refined the set by including only image types relevant to financial data analysis: Line Chart, Bar Chart / Column Chart, Diagram / Schematic, Candlestick Chart / OHLC Chart, and Infographic. Content themes were balanced across categories such as macroeconomic indicators, financial markets, risk management, and economic cycles to ensure thematic diversity.
After automated filtering, a manual review was conducted to exclude any remaining unsuitable images, such as those with poor readability, non-compliant elements, or ambiguous content. The final curated dataset consists of 3,092 images. The resulting dataset supports rigorous computational analysis of financial visualizations with high compliance and varying complexity.
Initial Images Statistics Summary. A total of 820,816 images were collected and annotated from a substantial collection of financial documents. The initial dataset’s distributions across content themes, image types, compliance levels, and complexity degrees are presented in Table 8.
Category Classification Count Percentage
Content Theme Macroeconomic Indicators 2.66%
Financial Markets & Products 63.20%
Commodities & Real Estate 0.99%
Bonds & Fixed Income 0.73%
Monetary & Fiscal Policy 0.88%
International Trade 0.88%
Corporate Finance & Valuation 1.39%
Industry Analysis 2.98%
Investment Theory & PM 13.42%
Risk Models & Management 1.13%
Economic Cycles & Theories 3.06%
Microeconomic Principles 0.43%
Demographics & Socioeconomics 3.57%
Financial Systems 1.17%
Organization & Regulation 1.35%
Geospatial Economic Data 1.62%
Financial History 0.53%
Total 100.00%
(a) Content Theme Distribution
Category Classification Count Percentage
Chart Type Line Chart 11.82%
Bar/Column Chart 5.52%
Pie/Donut Chart 0.43%
Scatter/Bubble Chart 0.45%
Table 2.50%
Diagram/Schematic 17.48%
Radar Chart 0.07%
Heatmap 0.05%
Candlestick/OHLC 46.43%
Photograph 3.99%
Infographic 11.27%
Total 100.00%
(b) Chart Type Distribution
Category Classification Count Percentage
Compliance Level 1 194,325 23.67%
2 3,703 0.45%
3 995 0.12%
4 9 0.00%
5 117 0.01%
6 614 0.07%
7 34,455 4.20%
8 561,368 68.39%
9 23,206 2.83%
10 2,024 0.25%
Total 820,816 100.00%
(c) Compliance Level Distribution
Category Classification Count Percentage
Complexity Degree 0 35 0.00%
1 52,841 6.44%
2 49,852 6.07%
3 49,937 6.08%
4 49,743 6.06%
5 82,867 10.10%
6 408,909 49.82%
7 126,258 15.38%
8 296 0.04%
9 78 0.01%
Total 820,816 100.00%
(d) Complexity Degree Distribution
Table 8: Initial Image Distributions
Sampling Images Statistics Summary. A total of 3,092 images were selected from the initial dataset for analysis. The distribution across image types and themes is visualized in Figure 1 (c) and (d), respectively. Regarding the compliance level of the final dataset, 2,642 images (85.45%) were rated at level 9, while 450 images (14.55%) were rated at level 10. For complexity degree, the majority of images, 2,741 (88.65%), were rated at level 7, followed by 188 (6.08%) at level 8, 140 (4.53%) at level 6, 17 (0.55%) at level 9, and 6 (0.19%) at level 5.
Appendix C Limitations
We identify several key limitations in our current work:
Potential Data Contamination. Since PyFi-600K is derived from publicly available financial reports and books, there is a risk that these documents were included in the pre-training corpora of the evaluated VLMs. To address this, we implemented rigorous data-level filtering. We conducted "blind tests" to determine if questions could be answered correctly without visual input and "completion tests" to check if models could reconstruct content from text alone. Samples where models succeeded without visual context were flagged as potential leakage and excluded from our core evaluation set to ensure that performance gains reflect genuine multimodal reasoning. Although it may always be beneficial to include documents published after the release of large language models, doing so can significantly reduce the scalability and size of the dataset, introduce manual effort, and prevent the data from being updated continuously.
Computational Constraints on Scale. Although PyFi-600K contains 600k samples, computational limits restricted us from scaling further. However, the pipeline is designed to be extensible and can leverage the vast array of financial datasets available (e.g., on Hugging Face) to construct billion-scale datasets in the future.
Reward Modeling Noise. The reward scores used during the adversarial interactions and MCTS may contain noise. Accurately evaluating open-ended financial reasoning is inherently challenging, and minor inaccuracies in the reward model can affect the optimization process.
Adversarial Framework Complexity. The PyFi-adv mechanism, while effective for data synthesis, introduces complexity that may limit its deployment in resource-constrained environments. Furthermore, the stochastic nature of VLMs can occasionally lead to inconsistent chain lengths or hallucinations, which requires careful post-hoc filtering.
Incomplete Process Supervision Analysis. Due to time and resource limitations, we have not yet fully explored the upper bounds of process supervision provided by our framework, although our initial experiments demonstrate its significant value.
Appendix D Discussion on Data Leakage
In constructing PyFi-600K, we acknowledge the potential risk of data leakage, particularly due to the use of publicly available financial documents that may have been included in the pretraining corpora of some VLMs. To ensure a fair and reliable evaluation, we conduct a systematic and stringent analysis to identify and mitigate such leakage.
Existing research employs various methods to detect data leakage. Dong et al. (2024) identify leakage by measuring the peakedness of the model’s output distribution, where high similarity among sampled outputs indicates memorization. Wu et al. (2025b) assess leakage by the model’s ability to reconstruct or answer problems based only on partial prompt prefixes. Ni et al. (2025) detect leakage in multiple-choice benchmarks by identifying if the log-probability of the original option order is a significant outlier among permutations. However, these approaches are often indirect, model-specific, or limited to particular benchmark formats, making them less applicable to financial image understanding tasks where the interplay between image and text is essential.
We define a sample as exhibiting data leakage if more than half of the evaluated models can answer the question correctly without access to the image. This criterion is based on the assumption that if models can consistently answer a question without visual input, the answer is likely encoded in the textual content or has been memorized during pretraining, rather than being inferred from the image. Our method is more direct, task-agnostic, and interpretable: by explicitly ablating the visual modality and measuring performance across diverse models, we capture leakage that genuinely compromises multimodal evaluation fairness—making it the most suitable strategy for benchmarks like PyFi-600K that hinge on visual grounding.
To quantify this, we conduct a comprehensive study where we evaluate all 15 VLMs on the original 1,000-sample test set under two conditions: with and without the image input. Samples that meet the leakage criterion are subsequently excluded from the evaluation set. As a result, the test set is reduced from 1,000 to 301 samples, forming the Evaluation Dataset after Exclusion. All results reported in the main paper are derived exclusively from this rigorously filtered set.
This filtered dataset is used for all final evaluations and fine-tuning experiments reported in the main paper. The significant reduction in sample size underscores the prevalence of potential leakage in financial Q&A datasets and highlights the importance of this filtering step.
This approach is both reasonable and effective in mitigating the influence of data leakage. By removing samples that can be answered without visual reasoning, we ensure that the remaining questions require genuine visual understanding and financial reasoning, aligning with the goals of PyFi-600K. Furthermore, this method is model-agnostic and does not rely on external knowledge of pretraining data, making it broadly applicable.
In summary, our data leakage mitigation strategy enhances the validity of the evaluation results and supports the development of more robust and generalizable financial VLMs.
Appendix E Discussion on Pyramid-like Structure
Financial image analysis requires more than object detection or simple classification; true understanding is a multi-stage cognitive process from visual perception to strategic decision-making. To address this, we propose PyFi-600K, a pyramid-like hierarchy structured across six levels: Perception, Data Extraction, Calculation Analysis, Pattern Recognition, Logical Reasoning, and Decision Support.
Level 1: Perception. This is the foundational base of the pyramid. At this stage, the model performs basic visual recognition. It identifies fundamental elements within the image: "What objects are present?" or "What is the color of that line?" This involves detecting axes, labels, data points, lines, bars, pie segments, and textual annotations. The questions and answers are a structured inventory of visual components.
Level 2: Data Extraction. Building on the perceived components, this level translates visual information into machine-readable quantitative or categorical data. For example, it converts the height of a bar in a chart into a numerical value based on the Y-axis scale, or it extracts the exact numerical percentage from a pie chart label. The output is a clean dataset ready for computation.
Level 3: Calculation Analysis. Here, the extracted data undergoes initial processing. This involves performing basic arithmetic, such as calculating percentage changes, growth rates or variances, and descriptive statistics. This level answers "What are the basic quantitative relationships?"
Level 4: Pattern Recognition. This level moves from discrete calculations to identifying broader trends and shapes within the data. It answers "What is happening over time or across categories?" Examples include recognizing an upward or downward trend in a stock price chart, identifying a period of high volatility, or spotting seasonal patterns in sales data.
Level 5: Logical Reasoning. This is a critical step towards higher-order cognition. The model integrates the identified patterns with external or common-sense financial knowledge to answer "Why is this happening, or what does this imply?" For instance, reasoning that a sharp drop in a company’s stock price coinciding with a negative earnings report indicates a causal relationship. It involves drawing inferences and establishing cause-and-effect links.
Level 6: Decision Support. The apex of the pyramid leverages all previous levels to provide actionable insights. It answers "What should be done?" Based on the reasoned analysis and the options or answers might include recommendations.
PyFi-600K provides a robust and principled framework for building a multi-modal financial image understanding dataset. By decomposing the task into six hierarchically organized cognitive levels for automated, reward-driven generation of logical chains, we create a resource that promotes deep, structured, and explainable reasoning.
Appendix F Discussion on Question Chain
A complete question chain can refer to the Figure 12, Figure 13, and Figure 14.
Figure 12: Illustration of the complete question chain presented in the Figure 2 of the main paper.
Complete Example of Question Chain 1. Question: Which shaded region in the chart represents the ’Period of low-oil prices’?
Options:
 A. Gray (Phase 1: pre-oil period)
 B. Purple (Phase 2: post-oil period)
 C. Light blue (Phase 3: Period of low-oil prices)
 D. Green (not a shaded region)
Answer: C
2. Question: What is the color of the shaded region labeled ’Phase 2: post-oil period’ in the chart?
Options:
 A. Light gray
 B. Dark gray
 C. Light blue
 D. Cyan
Answer: B
3. Question: During Phase 3: Period of low-oil prices, what is the approximate value of general government revenue (blue dashed line) in Chad, as shown in the chart?
Options:
 A. 5% of GDP
 B. -2.4% of GDP
 C. 10% of GDP
 D. 20% of GDP
Answer: B
4. Question: Which event is associated with the decline in general government revenue (blue dashed line) during Phase 3 (Period of low-oil prices) in Chad, as shown in the chart?
Options:
 A. Internal conflict
 B. Drought
 C. Drop in oil prices
 D. Normalisation of ties with Sudan
Answer: C
5. Question: In 2016, during the Period of low-oil prices (Phase 3), what is the approximate value of (general government total expenditure - general government revenue) in Chad?
Options:
 A.  17.4%
 B.  -17.4%
 C.  5%
 D.  -5%
Answer: A
6. Question: What pattern is observed between general government revenue and total expenditure during the ’Onset oil production’ period (marked in the chart)?
Options:
 A. Both revenue and expenditure increase significantly
 B. Revenue increases while expenditure decreases
 C. Revenue decreases while expenditure increases
 D. Both revenue and expenditure remain stable
Answer: A
Figure 13: Complete example of question chain.
Complete Example of Question Chain (continued) 7. Question: What pattern does the general government total expenditure (green dashed line) follow during periods of internal conflict (marked by red arrows) in Chad from 1990 to 2017?
Options:
 A. It decreases consistently
 B. It increases significantly
 C. It remains stable
 D. It fluctuates randomly
Answer: B
8. Question: How does the volatility of general government revenue in Chad during the Period of low-oil prices (Phase 3) contribute to economic vulnerability?
Options:
 A. It increases military spending, leading to internal conflict.
 B. It reduces the government’s ability to withstand shocks and invest in stability.
 C. It causes oil prices to drop further, worsening revenue.
 D. It eliminates the need for non-oil revenue sources.
Answer: B
9. Question: Based on the image illustrating Chad’s economic volatility driven by oil price cycles and security shocks, how would establishing a sovereign wealth fund help reduce revenue vulnerability during low-oil price periods?
Options:
 A. By increasing military spending to preempt conflict and stabilize revenue
 B. By investing oil revenues during high-price periods to fund expenditures during low-price periods
 C. By reducing reliance on non-oil revenue sources, such as IMF-ECF support
 D. By increasing general government revenue through borrowing from the World Bank during booms
Answer: B
10. Question: Which policy would most effectively address both Chad’s oil revenue volatility (from external shocks) and vulnerability to internal conflict?
Options:
 A. Increase military spending to preempt conflict
 B. Establish a sovereign wealth fund to stabilize revenues and investments
 C. Reduce all non-oil revenue sources to avoid over-reliance on oil
 D. Borrow heavily from the World Bank during low-oil price periods
Answer: B
11. Final Question: Based on the image illustrating Chad’s fiscal volatility driven by oil cycles and security shocks, which policy would most effectively reduce economic vulnerability to future external shocks and internal conflict?
Options:
 A. Increase military spending to preempt conflict
 B. Establish a sovereign wealth fund to stabilize revenues and investments
 C. Reduce all non-oil revenue sources to avoid over-reliance on oil
 D. Borrow heavily from the World Bank during low-oil price periods
Answer: B
Figure 14: Complete example of question chain (continued).
Appendix G Model Performance with and without Fine-Tuning
The examples of Qwen2.5-VL-3B-Instruct performance before and after fine-tuning can refer to the Figures 15, 16, and 17, and the examples of Qwen2.5-VL-7B-Instruct can refer to the Figures 18, 19, 20 and 21.
Figure 15: Model performance before and after fine-tuning (Qwen2.5-VL-3B-Instruct)
Figure 16: Model performance before and after fine-tuning (Qwen2.5-VL-3B-Instruct)
Figure 17: Model performance before and after fine-tuning (Qwen2.5-VL-3B-Instruct)
Figure 18: Model performance before and after fine-tuning (Qwen2.5-VL-7B-Instruct)
Figure 19: Model performance before and after fine-tuning (Qwen2.5-VL-7B-Instruct)
Figure 20: Model performance before and after fine-tuning (Qwen2.5-VL-7B-Instruct)
Figure 21: Model performance before and after fine-tuning (Qwen2.5-VL-7B-Instruct)
Appendix H Detailed Results
H.1 Experiments Consideration and Ablation Studies
In order to comprehensively evaluate the performance of various VLMs on our samples from PyFi-600K and eliminate potential biases, we consider the following factors and conduct comparative experiments and ablation studies accordingly.
The Impact of Contextual Information. To investigate the impact of image-related contextual information on model performance, we provide the target model with an image and a question, followed by additional textual information in one of four conditions: (1) providing image background information only, (2) providing image contextual analysis only, (3) providing both background information and contextual analysis, and (4) providing no additional text. The background information includes the image’s title, topic, and the full forms of all abbreviations, offering necessary context for answering the question without revealing the answer itself. The contextual analysis, generated by a large language model, summarizes relevant contextual information present in the image and offers direct guidance for solving the problem. This experimental design aims to determine whether the model genuinely understands the visual content or relies on external textual information to infer the answer.
Risk of Data Leakage. The VLMs evaluated in this study were pretrained on massive corpora. Since our data primarily consists of publicly available financial documents sourced from the internet and predominantly dated before 2025, there is a potential risk of data leakage. To address this concern, we additionally evaluated the VLMs in a setting where images were not provided. By comparing model accuracy between the two settings, with and without images, we can assess whether the pretraining corpus might already contain images or questions from our dataset. As for the formal statistics, we exclude those samples that can be correctly answered by more than half of the evaluated models even without image input, and the remaining samples are defined as Evaluation Dataset after Exclusion while the initial ones are defined as Original Evaluation Dataset.
Impact of Option Distribution. Our analysis of the ground-truth labels reveals severe class imbalance. Specifically, 43.39% of the samples designate option A as the correct answer, 34.20% select option B, 15.81% choose option C, and only 6.08% pick option D. All remaining options combined account for less than 1% of the total answers. Details can be found in the appendix 7. Such extreme imbalance may introduce bias during both training and evaluation, potentially undermining model performance. To address this issue and obtain more robust and reliable evaluation results, we rebalance the sample options so that, in a set of one thousand samples, each of the four options A, B, C, and D constitutes approximately 25%. We then compare model performance before and after this rebalancing.
All experimental results are presented in the following sections. The meanings of key abbreviations and the group configurations are defined in the tables below.
Table 9: Meaning of Abbreviations
Abbreviation Meaning PP Perception DE Data Extraction CA Calculation Analysis PR Pattern Recognition LR Logical Reasoning DS Decision Support
Table 10: Experimental Group Configurations
Group Input Configuration I Image and Image Background II Image and Analysis Information III Image, Image Background and Analysis Information IV Image Only
H.2 Detailed Experiment Results on Original Evaluation Dataset with Image
Table 11: Model Performance Comparison By Group On Original Evaluation Dataset (Samples=1000, Groups with Image)
Model Overall TC TQ I II III IV Pre-trained VLMs GPT-4.1 76.08 3043 4000 75.4 75.0 77.5 76.4 InternVL3-38B 73.72 2949 4000 73.9 74.6 73.7 72.7 Claude-opus-4-1-20250805 79.88 3195 4000 80.8 79.5 80.6 78.6 GLM-4.5V 85.22 3409 4000 85.8 85.3 85.5 84.3 DeepSeek-VL2 63.20 2528 4000 61.8 63.3 65.6 62.1 Hunyuan-Large-Vision 76.60 3064 4000 77.3 76.4 77.5 75.2 ERNIE-4.5-turbo-vl 68.83 2753 4000 68.0 70.0 71.1 66.2 Moonshot-V1-128k-Vision-Preview 76.40 3056 4000 75.3 77.4 78.2 74.7 Moonshot-V1-32k-Vision-Preview 76.22 3049 4000 75.1 77.1 77.8 74.9 Moonshot-V1-8k-Vision-Preview 76.53 3061 4000 75.0 77.5 77.8 75.8 Qwen3-VL-Plus 76.70 3068 4000 76.5 76.6 76.8 76.9 Qwen2.5-VL-72B-Instruct 74.28 2971 4000 72.9 75.4 75.5 73.3 Qwen2.5-VL-32B-Instruct 71.78 2871 4000 71.2 72.4 73.0 70.5 Qwen2.5-VL-7B-Instruct 67.55 2702 4000 65.0 68.4 69.7 67.1 Qwen2.5-VL-3B-Instruct 59.00 2360 4000 57.3 59.7 62.1 56.9 SFT VLMs, Trained with 500 Sample Chains Qwen2.5-VL-3B-Instruct-L 58.80 2352 4000 58.4 60.3 60.0 56.5 PyFi-QwenVL-3B-500 61.52 2461 4000 60.6 62.8 63.6 59.1 PyFi-QwenVL-3B-COT-500 62.32 2493 4000 61.8 63.9 63.3 60.3 Qwen2.5-VL-7B-Instruct-L 66.60 2664 4000 65.5 67.5 68.4 65.0 PyFi-QwenVL-7B-500 66.27 2651 4000 62.3 67.1 67.2 68.5 PyFi-QwenVL-7B-COT-500 65.72 2629 4000 64.4 66.4 66.0 66.1 SFT VLMs, Trained with 47K Sample Chains Qwen2.5-VL-3B-Instruct-L 58.80 2352 4000 58.4 60.3 60.0 56.5 PyFi-QwenVL-3B-47K 61.85 2474 4000 61.6 63.0 62.5 60.3 PyFi-QwenVL-3B-COT-47K 64.40 2576 4000 64.2 63.6 66.6 63.2 Qwen2.5-VL-7B-Instruct-L 66.60 2664 4000 65.5 67.5 68.4 65.0 PyFi-QwenVL-7B-47K 66.35 2654 4000 65.7 67.1 67.2 65.4 PyFi-QwenVL-7B-COT-47K 67.42 2697 4000 67.8 67.6 68.0 66.3
Table 11 presents the overall accuracy and performance across groups for all models on the original evaluation dataset (1000 samples). GLM-4.5V achieves the highest overall accuracy (85.22%) among all pre-trained models, significantly outperforming others. Claude-opus follows as the second-best pre-trained model (79.88%). Among the Qwen series, the larger models generally perform better, with Qwen3-VL-Plus (76.70%) leading, showing a clear scaling effect from 3B (59.00%) to 72B (74.28%). For SFT models, fine-tuning with 500 or 47K samples shows modest but inconsistent improvements over their base Instruct-L versions. For instance, PyFi-QwenVL-3B-COT-47K reaches 64.40%, a noticeable gain over the base Qwen2.5-VL-3B-Instruct (59.00%). However, for the 7B models, gains from SFT are marginal, with some variants even underperforming the base model. This indicates that the effectiveness of SFT may be model-size and data-dependent.
The following tables present performance on the Evaluation Dataset after Exclusion, which removes samples that more than half the models could answer correctly without the image. This subset (301 samples) is designed to mitigate data leakage concerns and focus on problems that genuinely require visual understanding.
Table 12: Model Performance Comparison By Group On Evaluation Dataset After Exclusion (Samples=301, Groups with Image)
Model Overall TC TQ I II III IV Pre-trained VLMs GPT-4.1 52.99 638 1204 51.5 52.16 52.82 55.48 InternVL3-38B 52.91 637 1204 55.81 52.16 50.17 53.49 Claude-opus-4-1-20250805 64.70 779 1204 65.78 62.46 63.12 67.44 GLM-4.5V 74.75 900 1204 74.75 75.08 73.75 75.42 DeepSeek-VL2 45.18 544 1204 45.85 42.86 45.18 46.84 Hunyuan-Large-Vision 59.72 719 1204 60.13 60.8 59.47 58.47 ERNIE-4.5-turbo-vl 34.47 415 1204 32.56 34.55 34.88 35.88 Moonshot-V1-128k-Vision-Preview 54.57 657 1204 53.82 54.15 54.15 56.15 Moonshot-V1-32k-Vision-Preview 54.40 655 1204 54.82 53.82 53.16 55.81 Moonshot-V1-8k-Vision-Preview 54.90 661 1204 53.82 54.15 54.15 57.48 Qwen3-VL-Plus 51.00 614 1204 52.16 49.17 48.17 54.49 Qwen2.5-VL-72B-Instruct 48.84 588 1204 47.18 47.51 46.51 54.15 Qwen2.5-VL-32B-Instruct 43.19 520 1204 40.86 42.19 42.19 47.51 Qwen2.5-VL-7B-Instruct 37.87 456 1204 36.21 36.21 36.54 42.52 Qwen2.5-VL-3B-Instruct 20.51 247 1204 18.6 18.94 21.26 23.26 SFT VLMs, Trained with 500 Sample Chains Qwen2.5-VL-3B-Instruct-L 20.85 251 1204 20.6 21.93 17.28 23.59 PyFi-QwenVL-3B-500 26.74 322 1204 28.57 25.25 23.92 29.24 PyFi-QwenVL-3B-COT-500 27.66 333 1204 29.57 25.58 21.59 33.89 Qwen2.5-VL-7B-Instruct-L 34.55 416 1204 34.55 32.89 32.23 38.54 PyFi-QwenVL-7B-500 43.44 523 1204 39.53 41.53 40.2 52.49 PyFi-QwenVL-7B-COT-500 39.53 476 1204 40.86 38.87 34.88 43.52 SFT VLMs, Trained with 47K Sample Chains Qwen2.5-VL-3B-Instruct-L 20.85 251 1204 20.6 21.93 17.28 23.59 PyFi-QwenVL-3B-47K 25.25 304 1204 25.58 24.25 22.92 28.24 PyFi-QwenVL-3B-COT-47K 40.37 486 1204 40.2 37.21 38.87 45.18 Qwen2.5-VL-7B-Instruct-L 34.55 416 1204 34.55 32.89 32.23 38.54 PyFi-QwenVL-7B-47K 27.08 326 1204 26.58 25.91 24.92 30.9 PyFi-QwenVL-7B-COT-47K 42.61 513 1204 44.19 38.87 42.52 44.85
Table 12 provides an overall summary on the excluded dataset, including total correct (TC) out of total questions (TQ = 4 groups * 301 samples = 1204). The ranking of pre-trained models remains consistent. It also shows performance per group (I, II, III, IV). The top models like GLM-4.5V and Claude-opus maintain high and consistent accuracy across all four groups, demonstrating robustness. The SFT results reinforce previous observations: The 47K COT training for the 3B model (PyFi-QwenVL-3B-COT-47K) yields the most dramatic improvement, boosting overall accuracy from 20.85% to 40.37% and showing gains across all groups. For the 7B models, the 500-sample SFT (PyFi-QwenVL-7B-500) provides a strong boost, but the 47K COT variant also delivers competitive performance. Notably, the non-COT 47K SFT for the 7B model (PyFi-QwenVL-7B-47K) underperforms significantly, highlighting the critical importance of the Chain-of-Thought training methodology alongside data volume for achieving gains on this challenging visual reasoning benchmark.
Table 13: Model Performance Comparison By Capability On Evaluation Dataset After Exclusion (Samples=301, Groups with Image)
Model Overall PP DE CA PR LR DS Pre-trained VLMs GPT-4.1 52.99 86.84 47.64 40.62 57.14 54.84 38.46 InternVL3-38B 52.91 76.32 56.37 44.53 43.88 53.23 30.77 Claude-opus-4-1-20250805 64.70 80.92 62.74 66.02 58.16 72.58 32.69 GLM-4.5V 74.75 89.47 78.07 65.23 75.51 75.81 46.15 DeepSeek-VL2 45.18 76.32 44.58 36.33 42.35 42.74 19.23 Hunyuan-Large-Vision 59.72 90.79 58.96 36.72 63.27 69.35 51.92 ERNIE-4.5-turbo-vl 34.47 34.87 33.96 36.72 25.00 45.16 36.54 Moonshot-V1-128k-Vision-Preview 54.57 80.92 61.08 39.06 41.84 58.87 38.46 Moonshot-V1-32k-Vision-Preview 54.40 80.26 61.56 37.11 40.31 62.90 38.46 Moonshot-V1-8k-Vision-Preview 54.90 80.26 61.79 39.84 42.35 58.06 38.46 Qwen3-VL-Plus 51.00 77.63 54.25 37.50 41.84 57.26 32.69 Qwen2.5-VL-72B-Instruct 48.84 70.39 59.91 33.20 35.71 46.77 26.92 Qwen2.5-VL-32B-Instruct 43.19 60.53 45.28 37.50 36.73 42.74 28.85 Qwen2.5-VL-7B-Instruct 37.87 54.61 49.53 34.77 18.37 24.19 15.38 Qwen2.5-VL-3B-Instruct 20.51 36.84 15.57 21.48 16.84 21.77 19.23 SFT VLMs, Trained with 500 Sample Chains Qwen2.5-VL-3B-Instruct-L 20.85 35.53 17.22 21.09 15.82 20.97 25.00 PyFi-QwenVL-3B-500 26.74 34.87 28.77 28.52 16.84 24.19 21.15 PyFi-QwenVL-3B-COT-500 27.66 37.50 28.77 26.17 22.45 25.81 21.15 Qwen2.5-VL-7B-Instruct-L 34.55 52.63 42.92 31.64 18.37 25.00 11.54 PyFi-QwenVL-7B-500 43.44 71.05 44.58 43.36 30.10 37.90 17.31 PyFi-QwenVL-7B-COT-500 39.53 61.84 41.51 40.23 26.02 34.68 17.31 SFT VLMs, Trained with 47K Sample Chains Qwen2.5-VL-3B-Instruct-L 20.85 35.53 17.22 21.09 15.82 20.97 25.00 PyFi-QwenVL-3B-47K 25.25 42.76 19.81 18.75 29.59 29.03 25.00 PyFi-QwenVL-3B-COT-47K 40.37 62.50 36.56 29.69 41.84 48.39 34.62 Qwen2.5-VL-7B-Instruct-L 34.55 52.63 42.92 31.64 18.37 25.00 11.54 PyFi-QwenVL-7B-47K 27.08 36.84 23.35 30.47 17.35 37.10 25.00 PyFi-QwenVL-7B-COT-47K 42.61 68.42 41.75 34.77 32.65 49.19 34.62
Table 13 breaks down model performance by required capability on the challenging 301-sample subset. The overall accuracy drops dramatically for all models compared to the full dataset, confirming the difficulty of these leakage-filtered samples. GLM-4.5V again dominates with 74.75% overall, excelling particularly in Perception (PP, 89.47%) and Data Extraction (DE, 78.07%). Claude-opus (64.70%) shows strong performance in Calculation Analysis (CA, 66.02%) and Logical Reasoning (LR, 72.58%). Hunyuan-Large-Vision is notable for its high Perception (90.79%) and relatively strong Decision Support (DS, 51.92%). A clear capability-wise difficulty trend emerges: Perception is generally easiest for top models, while Calculation Analysis and, especially, Decision Support are most challenging. Among SFT models, PyFi-QwenVL-3B-COT-47K shows a remarkable leap to 40.37%, more than doubling the performance of its base model (20.85%) and significantly outperforming its 500-sample counterpart. This suggests that Chain-of-Thought (COT) training on a larger, high-quality dataset (47K) is highly effective for smaller models, particularly improving Pattern Recognition (PR) and Logical Reasoning (LR). For the 7B models, PyFi-QwenVL-7B-500 achieves the best overall (43.44%), but the 47K COT variant also shows strong gains in LR.
Table 14: Model Performance Comparison By Complexity Degree On Evaluation Dataset After Exclusion (Samples=301, Groups with Image)
Model Overall 1 2 3 4 5 Pre-trained VLMs GPT-4.1 52.99 74.51 46.12 52.63 46.55 53.85 InternVL3-38B 52.91 75.00 52.91 41.78 44.83 67.31 Claude-opus-4-1-20250805 64.70 76.47 61.89 66.12 55.60 73.08 GLM-4.5V 74.75 87.75 76.21 72.37 64.22 73.08 DeepSeek-VL2 45.18 67.65 43.45 34.54 42.24 46.15 Hunyuan-Large-Vision 59.72 80.88 54.85 50.00 57.76 80.77 ERNIE-4.5-turbo-vl 34.47 32.84 33.25 38.82 25.43 65.38 Moonshot-V1-128k-Vision-Preview 54.57 74.02 56.31 46.71 43.97 57.69 Moonshot-V1-32k-Vision-Preview 54.40 73.53 56.31 45.72 43.97 61.54 Moonshot-V1-8k-Vision-Preview 54.90 73.53 56.80 48.03 43.97 55.77 Qwen3-VL-Plus 51.00 72.55 50.00 45.72 37.50 65.38 Qwen2.5-VL-72B-Instruct 48.84 67.16 57.77 36.51 31.03 57.69 Qwen2.5-VL-32B-Instruct 43.19 58.33 43.69 39.14 30.60 59.62 Qwen2.5-VL-7B-Instruct 37.87 55.39 44.90 31.91 17.67 38.46 Qwen2.5-VL-3B-Instruct 20.51 31.86 12.62 23.03 18.97 30.77 SFT VLMs, Trained with 500 Sample Chains Qwen2.5-VL-3B-Instruct-L 20.85 33.33 13.11 21.38 21.12 28.85 PyFi-QwenVL-3B 26.74 40.69 20.87 28.29 20.69 36.54 PyFi-QwenVL-3B-COT 27.66 39.22 22.57 27.96 24.57 34.62 Qwen2.5-VL-7B-Instruct-L 34.55 55.88 37.86 30.26 15.52 34.62 PyFi-QwenVL-7B 43.44 66.18 44.17 39.47 28.45 38.46 PyFi-QwenVL-7B-COT 39.53 58.82 39.32 36.51 23.71 53.85 SFT VLMs, Trained with 47K Sample Chains Qwen2.5-VL-3B-Instruct-L 20.85 33.33 13.11 21.38 21.12 28.85 PyFi-QwenVL-3B 25.25 37.25 18.93 19.74 31.47 32.69 PyFi-QwenVL-3B-COT 40.37 59.80 32.77 36.51 40.09 48.08 Qwen2.5-VL-7B-Instruct-L 34.55 55.88 37.86 30.26 15.52 34.62 PyFi-QwenVL-7B 27.08 35.29 21.60 26.64 25.86 46.15 PyFi-QwenVL-7B-COT 42.61 62.25 38.35 39.80 32.76 59.62
Table 14 analyzes performance stratified by problem complexity degree (1=Least, 5=Most Complex). A strong negative correlation between model performance and complexity degree is evident for most models. For example, GLM-4.5V scores 87.75% on degree-1 problems but drops to 64.22% on degree-4. Interestingly, performance on degree-5 (the most complex) does not always follow this monotonic decline; for models like Claude-opus, Hunyuan, and some Qwen variants, degree-5 accuracy is higher than degree-4. This may indicate that the highest complexity problems, while difficult, sometimes involve different (e.g., more structured) reasoning that some models can handle. The SFT models show interesting patterns: PyFi-QwenVL-3B-COT-47K demonstrates substantial improvements across all complexity levels, most notably on degrees 4 and 5 (40.09% and 48.08%, up from  21% and  29% in the base model). This suggests that the COT training on extensive data particularly enhances the model’s ability to tackle high-complexity, multi-step reasoning problems. The PyFi-QwenVL-7B-COT-47K model also shows strong gains on degree-5 problems (59.62%).
Summary of Key Findings
Performance Gap. A significant performance gap exists between top-tier proprietary models (GLM-4.5V, Claude-opus) and open-source models, especially on the challenging leakage-filtered subset.
Scaling Laws. For the Qwen family, model size strongly correlates with performance on this financial visual QA task.
Effectiveness of SFT: Supervised Fine-Tuning can substantially improve smaller models (3B), particularly when using a large (47K), high-quality dataset with Chain-of-Thought reasoning steps. Gains for larger models (7B) are more modest and sensitive to the training recipe.
Capability & Complexity. Perception and Data Extraction are relative strengths for leading models, while Calculation Analysis and Decision Support remain challenging. Problem complexity is inversely correlated with accuracy, but advanced training (COT) can specifically improve performance on high-complexity problems.
Appendix I Algorithms
The proposed adversarial agent dataset construction method is outlined in Algorithm 1. Its runtime is primarily dictated by the base VLM’s inference latency. Our implementation, using Doubao-seed-1-6-flash-250828, completes approximately 100 adversarial rounds (100 samples) in about 60 minutes. While the MCTS process for a single image runs sequentially, each adversarial step can generate multiple branches in parallel for internal acceleration. The framework also enables concurrent processing of multiple images across servers and processes. Consequently, the approach is readily extensible to large-scale synthesis with manageable time costs.
Input: Challenger Agent $\psi$, Solver Agent $\phi$, Image II with background $b_{I}$ and analysis info $f_{I}$, capability levels $\mathcal{L}=\{l_{1},\dots,l_{6}\}$, complexity levels $\mathcal{C}=\{c_{1},\dots,c_{5}\}$, MCTS parameters $\alpha,\beta$, max chains $T_{\max}$
Output: Pyramid-structured dataset $\mathcal{S}$ with question-answer chains
1
2Generate final question $Q_{F}$ using LLMs with: $b_{I}$, $f_{I}$, $\text{$l$}(Q_{F})=l_{6}$,
3 Determine ground truth answer $A_{F}$ for $Q_{F}$ via multi-LLM consensus
4
5Initialize MCTS tree $\mathcal{T}$ with root $Q_{F}$
6 $t\leftarrow 0$
7 while $t<T_{\max}$ and $\mathcal{T}$ not terminated do
8 ${\psi},{\phi}\leftarrow$ root of $\mathcal{T}$
9 while $\psi$} and $\phi$} ready and debate not terminated do
10    Compete (Challenger Agent $\psi$ Turn)
11    $N_{\text{child}}\leftarrow$ children of $q_{\psi}$
12    if $|N_{\text{child}}|=0$
13         Determine current and next capability level $l_{\text{curr}}$, $l_{\text{next}}$ based on pyramid progression
14         Generate new question $q_{\text{new}}$ with: $b_{I}$, $l_{\text{curr}}=l_{\text{next}}$, $\text{complexity}\in\mathcal{C}$, $\text{visit\_count}=1$, $\text{victory\_count}=0$
15       $q_{\psi}\leftarrow q_{\text{new}}$
16
17    else
18       $p_{\text{gen}}\leftarrow\sigma(\alpha\cdot(|N_{\text{child}}|-\beta))$
19         Sample $z\sim\text{Bernoulli}(p_{\text{gen}})$
20       if $z=1$
21            Determine current and next capability level $l_{\text{curr}}$, $l_{\text{next}}$ based on pyramid progression
22            Generate new question $q_{\text{new}}$ with: $b_{I}$, $\text{capability}=l_{\text{next}}$, $\text{complexity}\in\mathcal{C}$, $\text{visit\_count}=1$, $\text{victory\_count}=0$
23          $q_{\psi}\leftarrow q_{\text{new}}$
24
25       else
26          $q_{\psi}\leftarrow\arg\max_{n^{\prime}\in N_{\text{child}}}\text{UCT}(n^{\prime})$
27            Increment $\text{visit\_count}(q_{\psi})$
28
29         end if
30
31      end if
32
33    Compete (Solver Agent $\phi$ Turn)
34      Generate answer $a_{\phi}$ to $q_{\psi}$
35    if $a_{\phi}\in$ existing actions
36         Increment $\text{visit\_count}(a_{\phi})$
37
38    else
39         Create new $a_{\phi}$ with $\text{visit\_count}\leftarrow 1$, $\text{victory\_count}\leftarrow 0$
40
41      end if
42     Add edge $(q_{\psi},a_{\phi})$ to $\mathcal{T}$
43
44    if $l_{\text{curr}}$ = 6 and LLM judges chain can answer $Q_{F}$
45         Generate final answer $a_{\text{final}}$
46       if $a_{\text{final}}=A_{F}$
47            Backpropagate and increment victory_count for all nodes
48
49         end if
50        Break
51
52      end if
53
54   end while
55
56 $t\leftarrow t+1$
57
58 end while
59Add successful chains to $\mathcal{S}$
60
Algorithm 1 Dataset Construction via Adversarial Agents under the MCTS Paradigm